### v1

# Asymptotic regularity of a generalised stochastic Halpern scheme with applications

Abstract.
We provide abstract, general and highly uniform rates of asymptotic regularity for a generalized stochastic Halpern-style iteration, which incorporates a second mapping in the style of a Krasnoselskii-Mann iteration. This iteration is general in two ways: First, it incorporates stochasticity in a completely abstract way rather than fixing a sampling method; secondly, it includes as special cases stochastic versions of various schemes from the optimization literature, including Halpern’s iteration as well as a Krasnoselskii-Mann iteration with Tikhonov regularization terms in the sense of Boţ, Csetnek and Meier. For these particular cases, we in particular obtain linear rates of asymptotic regularity, matching (or improving) the currently best known rates for these iterations in stochastic optimization, and quadratic rates of asymptotic regularity are obtained in the context of inner product spaces for the general iteration. We utilize these rates to give bounds on the oracle complexity of such iterations under suitable variance assumptions and batching strategies, again presented in an abstract style. Finally, we sketch how the schemes presented here can be instantiated in the context of reinforcement learning to yield novel methods for Q-learning.
a Department of Mathematics, Technische Universität Darmstadt,
Schlossgartenstraße 7, 64289 Darmstadt, Germany,
b Department of Computer Science, University of Bath,
Claverton Down, Bath, BA2 7AY, United Kingdom,
E-mails: pischke@mathematik.tu-darmstadt.de, trjp20@bath.ac.uk
Keywords: Asymptotic regularity, Halpern iteration, Tikhonov regularization, Q-learning, proof mining
MSC2020 Classification: 47J25, 47H09, 62L20, 03F10
1. Introduction
Approximating fixed points of nonexpansive mappings is one of the most fundamental tasks in nonlinear analysis and optimization. The problem becomes particularly interesting when we only have noisy versions those mappings, in which case the resulting approximation methods become stochastic processes. Concrete examples of this general situation include model-free reinforcement learning algorithms, where variants of $Q$-learning, for instance, can be viewed as stochastic methods for computing fixpoints of nonexpansive operators.
To be more concrete, let $(X,\left\lVert\cdot\right\rVert)$ be a seperable real-valued normed space and $T,U:X\to X$ be two nonexpansive mappings on $X$, i.e.
$\left\lVert Tx-Ty\right\rVert\leq\left\lVert x-y\right\rVert\text{ and }\left\lVert Ux-Uy\right\rVert\leq\left\lVert x-y\right\rVert$
for all $x,y\in X$. In order to approximate common fixed points of two such mappings under stochastic noise constraints, we introduce in this paper the so-called stochastic Halpern-Mann iteration, given by the schema
$\begin{cases}y_{n}:=(1-\alpha_{n})(Tx_{n}+\xi_{n})+\alpha_{n}u,\\ x_{n+1}:=(1-\beta_{n})(Uy_{n}+\delta_{n})+\beta_{n}y_{n},\end{cases}$
where, over some fixed probability space $(\Omega,\mathcal{F},\mathbb{P})$, $x_{0}$ and $u$ are arbitrary $X$-valued random variables chosen as a fixed starting point and as an anchor of the iteration, respectively, $(\xi_{n}),(\delta_{n})$ are sequences of $X$-valued random variables representing the stochastic noise, and $(\alpha_{n}),(\beta_{n})\subseteq[0,1]$ are suitable nonstochastic parameter sequences.
The schema (1) is designed to capture situations in which one does not have direct access to $Tx_{n}$ (respectively $Uy_{n}$), but can only use noisy versions $\tilde{T}$ of $T$ and $\tilde{U}$ of $U$. Intuitively, $\xi_{n}$ would then represent the difference between $Tx_{n}$ and the corresponding approximation of $Tx_{n}$ obtained from $\tilde{T}$ through a suitable sampling method (and similarly for $\delta_{n}$ and $Uy_{n}$), though our presentation is fully abstract and we will make no assumptions about $\xi_{n}$ or $\delta_{n}$, other than imposing controls on the means $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$ and $\mathbb{E}[\left\lVert\delta\right\rVert_{n}]$, which however are carefully chosen so that they can be achieved in concrete scenarios through sampling methods such as minibatching under suitable assumptions.
The main results of the paper establish general conditions under which we guarantee the asymptotic regularity of the scheme (1), both in the traditional sense of $\left\lVert x_{n}-x_{n+1}\right\rVert$ (sometimes called the discrete velocity [2]) and also relative to the mappings, i.e. considering the displacements $\left\lVert x_{n}-Tx_{n}\right\rVert$ and $\left\lVert x_{n}-Ux_{n}\right\rVert$. Furthermore, we establish these asymptotic regularity results both in expectation and almost surely, i.e. we show both
$\mathbb{E}[\left\lVert x_{n}-x_{n+1}\right\rVert],\ \mathbb{E}[\left\lVert x_{n}-Tx_{n}\right\rVert],\ \mathbb{E}[\left\lVert x_{n}-Ux_{n}\right\rVert]\to 0$
as well as
$\left\lVert x_{n}-x_{n+1}\right\rVert,\ \left\lVert x_{n}-Tx_{n}\right\rVert,\ \left\lVert x_{n}-Ux_{n}\right\rVert\to 0\text{ almost surely}$
under suitable conditions on the auxiliary parameters as well as the noise. In the general case this requires a subtle geometric argument and involves a delicate interplay between the two modes of stochastic convergence, which manifest, respectively, in the additional assumptions that $X$ is uniformly convex and that the sequence $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ is uniformly integrable. Most importantly, in all cases we provide explicit convergence rates for these expressions, under very general conditions and dependent only on a few moduli witnessing quantitative aspects of our main assumptions. In particular, we identify natural circumstances under which these rates are very fast, reaching up to linear speed in special cases.
To date, only the very simple instance of (1) corresponding to a stochastic variant of Halpern’s method [15] has been previously studied, where setting $U:=\mathrm{Id}$ and $\delta_{n}:=0$ gives rise to the iteration
$x_{n+1}=(1-\alpha_{n})(Tx_{n}+\xi_{n})+\alpha_{n}u,$
considered in [4] (for finite dimensional normed spaces), a scheme which in the Euclidean setting has recently received a great deal of attention in the context of stochastic monotone inclusion problems [5, 11, 30, 43]. In all cases, controlling the variance of the noise terms is crucial for convergence, and this is just one of several elements that makes the analysis of stochastic schemes such as (1) markedly different from that of their nonstochastic counterparts, some of the others being a focus on oracle complexity, and the relevance of stochastic methods to machine learning.
Given this increasing interest in stochastic variants of classic methods, the purpose of the present paper is to broaden their current scope and provide a collection of generalised convergence results in which all of the aforementioned features (variance reduction, oracle complexity, applications in machine learning) are presented in the abstract.
With this in mind, our main scheme (1) represents a stochastic analogue of the deterministic Halpern-Mann scheme recently introduced in [13], which integrates two of the most prominent methods for approximating fixed points for nonexpansive mappings, the Krasnoselskii-Mann method [23, 31] and Halpern’s method [15] (see also [40] and [41]), with the intended gain of combining the beneficial features of both, in particular the strong convergence of Halpern’s method even in infinite dimensional spaces.
While we focus on the special case of the stochastic Halpern iteration (1) at several points in the paper, where it forms a useful example and where in particular our results on asymptotic regularity in that case generalize and improve those presented in [4], our method (1) is certainly not limited to this special case, and in line with [13] encompasses stochastic variants of other well-known deterministic methods.
An important example of this is represented by setting $T:=\mathrm{Id}$ and $\xi_{n}:=0$, as well as $u:=0$ and $\gamma_{n}:=1-\alpha_{n}$, whereby we obtain a version of the Krasnoselskii-Mann iteration with Tikhonov regularization terms considered in [1, 42] that now incorporates stochastic noise
$x_{n+1}=(1-\beta_{n})(U(\gamma_{n}x_{n})+\delta_{n})+\beta_{n}(\gamma_{n}x_{n})$
[…]
While the special cases of (1) and (1) exhibit linear rates of asymptotic regularity under very mild (quantitative) assumptions on the space, parameters and noise, the general iteration (1) proves to be more complex as mentioned before, depending to a large degree on the geometry of the space: In the general case, we assume $X$ to be uniformly convex, and the rate then depends on a corresponding modulus witnessing this uniform convexity. Our main geometric argument is carried out in a pointwise manner, yielding almost sure convergence (with rates). Nevertheless, the passage back to convergence in mean requires a uniform integrability assumption (which as we show can actually be imposed on the noise terms), and our rates then also depend on a corresponding modulus of uniform integrability, which we define and use for the first time in this paper. However, fast rates also apply to the general scheme (1), where in the special case of a uniformly convex space of power type $p\geq 2$, the complexity of those rates is of order $p$, and so in particular we get quadratic rates of asymptotic regularity for inner product spaces.
While all of our rates are novel, we emphasise that the qualitative asymptotic regularity results for any iteration contained in (1) that goes beyond the Halpern iteration, are also to the best of our knowledge new to the literature.
All of the results obtained in this paper are motivated via the methodology of the proof mining program, a subfield of mathematical logic which combines an abstract approach to proofs in mainstream mathematics with the extraction of computational information, such as bounds or rates, from those proofs. We refer to the seminal monograph [19] for a comprehensive overview of both theoretical as well as applied aspects of this program, along with the survey [21] for an overview of more recent applications to nonlinear analysis. Proof mining has been widely applied in nonlinear analysis, and has found particular success in providing quantitative convergence results for Halpern’s iteration and its many variants, with notable instances ranging from initial rates of asymptotic regularity for Halpern’s iteration given by Leuştean [26] and the first analysis of Wittmann’s proof of the strong convergence of Halpern’s iteration given in the seminal paper of Kohlenbach [20], to the extensions of these results to nonlinear context such as $\mathrm{CAT}(0)$-spaces as in [22] (by a logical analysis of a corresponding convergence proof by Saejung [37]). They also include extensions of the Halpern iteration [38] for the modified Mann iteration introduced in [17] (and extended to nonlinear spaces in [10]) as well as the Krasnoselskii-Mann iteration with Tikhonov regularization terms and its extensions as in [6, 7, 8, 12] (with [6] of particular note, as linear rates of asymptotic regularity are there obtained for the first time in the context of applications of proof mining). In particular, the definition of the deterministic Halpern-Mann method given in [13] and its corresponding convergence proof were motivated by these logical considerations, as were the recent rates of asymptotic regularity given for this iteration in [29].
The present work departs from the aforementioned case studies in nonlinear analysis in that it incorporates, for the first time, stochasticity. In this way it forms part of a recent advance of proof mining into probability theory, which comprises both new developments in the logical foundations of probability theory due to first author and Neri [32], together with applied results on the quantitative aspects of stochastic processes by the second author and Neri [33, 34]. In particular, the present paper is one of the first applications of proof mining to stochastic optimization: It represents a particularly interesting case study in this respect, in that it does not readily follow from analogous quantitative results in the deterministic setting (such as in [13, 29]), but requires a substantial arsenal of new quantitative ideas for this stochastic setting, as will be introduced and discussed in detail throughout the paper. We stress that while this logical perspective was crucial in obtaining the present results, the paper does not rely on any notions from logic at all.
The main motivation of this paper is not so much its novelty within the proof mining program, but the real applicability of the main results in concrete areas. For this reason, our last section is dedicated to reinforcement learning, where a novel method of Q-learning with Tikhonov regularization terms is sketched, and the asymptotic regularity results from the main theoretical part of this paper are utilized to give very good oracle complexity estimates for a specific mini-batching procedure that matches those obtained for Halpern’s iteration in [4]. We also emphasise that potential applications of the stochastic Halpern-Mann method are not limited to this, and in particular the presence of the second mapping in the schema (1) allows for novel instantiations in style of double Q-learning [16] that will be explored, together with the Q-learning method with Tikhonov regularization terms mentioned before, in a subsequent paper in more detail. Another promising application is represented by stochastic versions of the splitting methods with Tikhonov regularization terms as studied in [1], but we leave this for future work, as we do the study of the convergence of the method (1), where we expect that strong convergence results hold for this method also in infinite dimensional spaces, analogously to the deterministic case.
2. Preliminaries and basic lemmas
[…]
Throughout the paper, we will be concerned with quantitative variants of various notions and we here now briefly the discuss the key definitions of the main quantitative notions used in the paper:
Given a non-negative sequence of reals $(a_{n})$, a rate of convergence for $a_{n}\to 0$ is a function $\varphi:(0,\infty)\to\mathbb{N}$ such that
$\forall\varepsilon>0\,\forall n\geq\varphi(\varepsilon)\left(a_{n}<\varepsilon\right).$
An invertible and decreasing rate $\varphi$ immediately implies the asymptotic estimate $a_{n}\leq\varphi^{-1}(n)$ for all $n\in\mathbb{N}$ and this immediately implies a complexity bound on the sequence in terms of the commonly used big O notation, namely $(a_{n})=O(\varphi^{-1}(n))$.
Now, given a nonnegative stochastic process $(X_{n})$, a rate of convergence for $X_{n}\to 0$ almost surely is a function $\Phi:(0,\infty)^{2}\to\mathbb{N}$ such that
$\forall\lambda,\varepsilon>0\left(\mathbb{P}\left(\exists n\geq\Phi(\lambda,\varepsilon)\,(X_{n}\geq\varepsilon)\right)<\lambda\right).$
We note that whenever $\Phi$ is a rate of convergence for $X_{n}\to 0$ almost surely, then for every $\varepsilon>0$, $\Phi(\varepsilon,\cdot)$ is a rate of convergence for $\mathbb{P}\left(\sup_{n\geq N}\,(X_{n}\geq\varepsilon)\right)\to 0$ in the nonstochastic sense.
Further, given a non-negative sequence of reals $(a_{n})$, we later want to quantitatively witness the convergence or divergence of the series over that sequence. For that, if $\sum_{n=0}^{\infty}a_{n}<\infty$, we say that a function $\chi:(0,\infty)\to\mathbb{N}$ is a rate of convergence for that sum if
$\forall\varepsilon>0\left(\sum_{n=\chi(\varepsilon)}^{\infty}a_{n}<\varepsilon\right).$
If $\sum_{n=0}^{\infty}a_{n}=\infty$, we say that a function $\theta:\mathbb{N}\times(0,\infty)\to\mathbb{N}$ is a rate of divergence for that sum if
$\forall b>0\,\forall k\in\mathbb{N}\left(\sum_{n=k}^{\theta(k,b)}a_{n}\geq b\right).$
Naturally, any such modulus satisfies $\theta(k,b)\geq k$ for any $k\in\mathbb{N}$ and $b>0$.
We now collect some of the basic abstract convergence results that our paper relies on. The most crucial of these, on the asymptotic behavior of sequences of reals satisfying certain recursive inequalities is the following due to Xu [41], often called Xu’s lemma:
Lemma 2.1 ([41]).
Suppose that $(s_{n}),(c_{n})\subseteq[0,\infty)$ as well as $(a_{n})\subseteq[0,1]$ and $(b_{n})\subseteq\mathbb{R}$ satisfy
$s_{n+1}\leq(1-a_{n})s_{n}+a_{n}b_{n}+c_{n}$
for all $n\in\mathbb{N}$ where $\sum_{n=0}^{\infty}a_{n}=\infty$, lim $\limsup b_{n}\leq 0$ and $\sum_{n=0}^{\infty}c_{n}<\infty$. Then $s_{n}\to 0$.
We will in particular rely on a quantitative rendering of an instance of Xu’s lemma which is represented by the following lemma. This result is contained in [22, 28] (up to the way the errors and the moduli are phrased) therefore for brevity we omit the proof.
Lemma 2.2 (essentially [22, 28]).
Suppose that $(s_{n}),(c_{n})\subseteq[0,\infty)$ and $(a_{n})\subseteq[0,1]$ satisfy
$s_{n+1}\leq(1-a_{n})s_{n}+c_{n}$
for all $n\in\mathbb{N}$, and furthermore, that $K>0$ is an upper bound on $(s_{n})$, $\theta$ is a rate of divergence for $\sum_{n=0}^{\infty}a_{n}=\infty$ and $\chi$ a rate of convergence for $\sum_{n=0}^{\infty}c_{n}<\infty$. Then $s_{n}\to 0$ with rate
$\varphi_{K,\theta,\chi}(\varepsilon):=\theta\left(\chi\left(\frac{\varepsilon}{2}\right),\ln\left(\frac{2K}{\varepsilon}\right)\right)+1.$
We now extend this lemma to a probabilistic variant. For that, we first consider the following result which allows us to transfer quantitative information from convergence in mean for “almost-monotone” sequences of random variables to rates of almost sure convergence.
Lemma 2.3.
Let $(X_{n})$, $(C_{n})$ be nonnegative stochastic processes satisfying
$X_{n+1}\leq X_{n}+C_{n}$
almost surely for all $n\in\mathbb{N}$ and suppose furthermore that
  1. (a)
    $\sum_{i=0}^{\infty}\mathbb{E}[C_{i}]<\infty$ with rate $\chi$,
  2. (b)
    $\mathbb{E}[X_{n}]\to 0$ with rate $\varphi$.
Then $X_{n}\to 0$ almost surely, and with rate
$\psi(\lambda,\varepsilon):=\max\left\{\varphi(\lambda\varepsilon/2),\chi(\lambda\varepsilon/2)\right\}.$
Proof.
We first note that for any $n\in\mathbb{N}$, we have $\mathbb{E}\left[\sum_{i=n}^{\infty}C_{i}\right]=\sum_{i=n}^{\infty}\mathbb{E}[C_{i}]$ by the monotone convergence theorem. Now define a stochastic process $(U_{n})$ by $U_{n}:=X_{n}+\sum_{i=n}^{\infty}C_{i}$. Then we have
$U_{n+1}=X_{n+1}+\sum_{i=n+1}^{\infty}C_{i}\leq X_{n}+C_{n}+\sum_{i=n+1}^{\infty}C_{i}\leq X_{n}+\sum_{i=n}^{\infty}C_{i}=U_{n}$
almost surely for any $n\in\mathbb{N}$, and therefore the events $(U_{n}\geq\varepsilon)$ are monotone decreasing in $n$. In particular, using Markov’s inequality, we get for any $N\in\mathbb{N}$:
$\mathbb{P}\left(\exists n\geq N(U_{n}\geq\varepsilon)\right)=\mathbb{P}\left(U_{N}\geq\varepsilon\right)$ $\leq\frac{\mathbb{E}[U_{N}]}{\varepsilon}=\frac{\mathbb{E}[X_{N}]+\sum_{i=N}^{\infty}\mathbb{E}\left[C_{i}\right]}{\varepsilon}.$
Therefore if $N=\psi(\lambda,\varepsilon)$, we have
$\mathbb{P}\left(\exists n\geq N(U_{n}\geq\varepsilon)\right)\leq\frac{\mathbb{E}[X_{N}]+\sum_{i=N}^{\infty}\mathbb{E}\left[C_{i}\right]}{\varepsilon}<\frac{(\lambda\varepsilon/2+\lambda\varepsilon/2)}{\varepsilon}=\lambda.$
The result follows by observing that $X_{n}\leq U_{n}$ holds almost surely for all $n\in\mathbb{N}$ and thus
$\mathbb{P}(\exists n\geq N\,(X_{n}\geq\varepsilon))\leq\mathbb{P}(\exists n\geq N\,(U_{n}\geq\varepsilon))<\lambda$
for any $N$, and in particular for the $N$ defined above. ∎
The above lemma now allows us to give a stochastic version of Lemma 2.2:
Lemma 2.4.
Suppose that $(X_{n})$, $(C_{n})$ are nonnegative stochastic processes satisfying
$X_{n+1}\leq(1-a_{n})X_{n}+C_{n}$
almost surely for all $n\in\mathbb{N}$. Furthermore, suppose that
  1. (a)
    $\mathbb{E}[X_{n}]\leq K$ for all $n\in\mathbb{N}$,
  2. (b)
    $\sum_{i=0}^{\infty}a_{i}=\infty$ with rate of divergence $\theta$,
  3. (c)
    $\sum_{i=0}^{\infty}\mathbb{E}[C_{i}]<\infty$ with rate of convergence $\chi$.
Then $\mathbb{E}[X_{n}]\to 0$ with rate $\varphi_{K,\theta,\chi}$ defined as in Lemma 2.2, i.e.
$\varphi_{K,\theta,\chi}(\varepsilon):=\theta\left(\chi\left(\frac{\varepsilon}{2}\right),\ln\left(\frac{2K}{\varepsilon}\right)\right)+1,$
and further $X_{n}\to 0$ almost surely with rate
$\psi_{K,\theta,\chi}(\lambda,\varepsilon):=\varphi_{K,\theta,\chi}\left(\frac{\lambda\varepsilon}{2}\right).$
Proof.
Taking expectations on both sides we have
$\mathbb{E}[X_{n+1}]\leq(1-a_{n})\mathbb{E}[X_{n}]+\mathbb{E}[C_{n}]$
for all $n\in\mathbb{N}$ and so the rate for $\mathbb{E}[X_{n}]\to 0$ follows by Lemma 2.2. For the rate for the almost sure convergence, observe that $\chi(\lambda\varepsilon/4)\leq\varphi_{K,\theta,\chi}(\lambda\varepsilon/2)$ as $\theta(k,b)\geq k$. Hence, one can proceed as in the proof of Lemma 2.3 to show that
$\mathbb{P}\left(\exists n\geq N(U_{n}\geq\varepsilon)\right)\leq\frac{\mathbb{E}[X_{N}]+\sum_{i=N}^{\infty}\mathbb{E}\left[C_{i}\right]}{\varepsilon}$
for any $N$ where, using $\chi(\lambda\varepsilon/4)\leq\varphi_{K,\theta,\chi}(\lambda\varepsilon/2)$, we then can conclude $\mathbb{P}\left(\exists n\geq N(U_{n}\geq\varepsilon)\right)<\lambda$ for $N=\varphi_{K,\theta,\chi}(\lambda\varepsilon/2)$. ∎
3. Quantitative asymptotic regularity for the generalized stochastic Halpern scheme
In this section we now outline our main theoretical results and derive rates of asymptotic regularity for the iterations generated by the generalized stochastic Halpern scheme.
3.1. Basic results and rates of asymptotic regularity
We begin with some fundamental recursive inequalities for the iterations generated by the iteration (1):
Lemma 3.1 (essentially [29]).
Let $(x_{n}),(y_{n})$ be the sequences generated by (1). Then the following recurrence relations hold pointwise everywhere for all $n\in\mathbb{N}$:
(1) $\left\lVert y_{n+1}-y_{n}\right\rVert$ $\leq(1-\alpha_{n+1})\left(\left\lVert x_{n+1}-x_{n}\right\rVert+\left\lVert\xi_{n+1}-\xi_{n}\right\rVert\right)+|\alpha_{n+1}-\alpha_{n}|\cdot\left\lVert Tx_{n}+\xi_{n}-u\right\rVert,$
(2) $\left\lVert x_{n+2}-x_{n+1}\right\rVert$ $\leq\left\lVert y_{n+1}-y_{n}\right\rVert+(1-\beta_{n+1})\left\lVert\delta_{n+1}-\delta_{n}\right\rVert+|\beta_{n+1}-\beta_{n}|\cdot\left\lVert Uy_{n}+\delta_{n}-y_{n}\right\rVert.$
Proof.
For (1) we observe that
$\left\lVert y_{n+1}-y_{n}\right\rVert=$ $\left\lVert(1-\alpha_{n+1})(Tx_{n+1}+\xi_{n+1})-(1-\alpha_{n})(Tx_{n}+\xi_{n})+(\alpha_{n+1}-\alpha_{n})u\right\rVert$
$\leq$ $(1-\alpha_{n+1})\left\lVert(Tx_{n+1}+\xi_{n+1})-(Tx_{n}+\xi_{n})\right\rVert$
$+\left\lVert(\alpha_{n}-\alpha_{n+1})(Tx_{n}+\xi_{n})-(\alpha_{n}-\alpha_{n+1})u\right\rVert$
$\leq$ $(1-\alpha_{n+1})\left(\left\lVert x_{n+1}-x_{n}\right\rVert+\left\lVert\xi_{n+1}-\xi_{n}\right\rVert\right)+|\alpha_{n+1}-\alpha_{n}|\cdot\left\lVert Tx_{n}+\xi_{n}-u\right\rVert$
where for the last inequality we use that $T$ is nonexpansive. Similarly for (2) we have
$\left\lVert x_{n+2}-x_{n+1}\right\rVert=$ $\left\lVert(1-\beta_{n+1})(Uy_{n+1}+\delta_{n+1})+\beta_{n+1}y_{n+1}-(1-\beta_{n})(Uy_{n}+\delta_{n})-\beta_{n}y_{n}\right\rVert$
$\leq$ $\left\lVert(1-\beta_{n+1})(Uy_{n+1}+\delta_{n+1})-(1-\beta_{n+1})(Uy_{n}+\delta_{n})+\beta_{n+1}(y_{n+1}-y_{n})\right\rVert$
$+\left\lVert(1-\beta_{n+1})(Uy_{n}+\delta_{n})-(1-\beta_{n})(Uy_{n}+\delta_{n})-(\beta_{n}-\beta_{n+1})y_{n}\right\rVert$
$\leq$ $(1-\beta_{n+1})\left(\left\lVert Uy_{n+1}-Uy_{n}\right\rVert+\left\lVert\delta_{n+1}-\delta_{n}\right\rVert\right)+\beta_{n+1}\left\lVert y_{n+1}-y_{n}\right\rVert$
$+|\beta_{n+1}-\beta_{n}|\cdot\left\lVert Uy_{n}+\delta_{n}-y_{n}\right\rVert$
$\leq$ $\left\lVert y_{n+1}-y_{n}\right\rVert+(1-\beta_{n+1})\left\lVert\delta_{n+1}-\delta_{n}\right\rVert+|\beta_{n+1}-\beta_{n}|\cdot\left\lVert Uy_{n}+\delta_{n}-y_{n}\right\rVert$
where again we use nonexpansivity of the operator in the last step. ∎
We now move to our first quantitative result which presents a rate of asymptotic regularity for the sequence $(x_{n})$, both in expectation and in probability. For that we introduce a first central assumption on the boundedness of the iteration (1) in expectation, as commonly made in the literature (see e.g. hypothesis $(\mathrm{H}_{1})$ in [4] of which this assumption here is a natural extension to the generalised iteration (1)):
(Hyp) There exists a $K_{0}\in\mathbb{N}^{*}$ such that for all $n\in\mathbb{N}$:
$\mathbb{E}[\left\lVert Tx_{n}-u\right\rVert],\ \mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert],\ \mathbb{E}[\left\lVert Uu-u\right\rVert],\ \mathbb{E}[\left\lVert Uy_{n}-u\right\rVert]\leq K_{0}<\infty$.
Throughout, if not stated otherwise, we will assume the existence of such a $K_{0}$.
In the context of the asymptotic regularity results that hold almost surely, we will sometimes need to make a slightly stronger assumption that the random variables are actually $L^{1}$-bounded in the following sense:
(Hyp) There exists a nonnegative random variable $Y$ with $K_{0}\geq\mathbb{E}[Y]$ for some $K_{0}\in\mathbb{N}^{*}$
and for all $n\in\mathbb{N}$: $\left\lVert Tx_{n}-u\right\rVert,\left\lVert Uy_{n}-y_{n}\right\rVert,\left\lVert Uu-u\right\rVert,\left\lVert Uy_{n}-u\right\rVert\leq Y$ almost surely.
Contrary to the above (Hyp), which will essentially always be tacitly assumed, we will always be very explicit about when we actually need to assume the above hypothesis (Hyp). It is to be noted that both hypotheses are guaranteed in the presence of a common fixed point of $T$ and $U$, as will be later discussed in more detail (see Lemma 3.9).
In any case, under the assumption (Hyp), we can immediately derive a bound on the expectation of the discrete velocity and utilize that to derive our first rate of asymptotic regularity:
Theorem 3.2.
Let $(x_{n}),(y_{n})$ be the sequences generated by (1). Suppose that $\sum_{n=0}^{\infty}\alpha_{n}=\infty$ with rate of divergence $\theta$, that
$\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n+1}-\xi_{n}\right\rVert],\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n+1}-\delta_{n}\right\rVert],\sum_{n=0}^{\infty}|\alpha_{n+1}-\alpha_{n}|,\sum_{n=0}^{\infty}|\beta_{n+1}-\beta_{n}|<\infty$
[…]
$\varphi_{K,\theta,\chi}(\varepsilon):=\theta\left(\chi\left(\frac{\varepsilon}{2}\right),\ln\left(\frac{2K}{\varepsilon}\right)\right)+1\text{ and }\psi_{K,\theta,\chi}(\lambda,\varepsilon):=\varphi_{K,\theta,\chi}\left(\frac{\lambda\varepsilon}{2}\right),$
where
$\chi(\varepsilon):=\max\{\chi_{1}(\varepsilon/4),\chi_{2}(\varepsilon/4),\chi_{3}(\varepsilon/4(E_{0}+K_{0}),\chi_{4}(\varepsilon/4(D_{0}+K_{0}))\}$
as well as $K:=2K_{0}+E_{0}+D_{0}+B$ for $B:=B_{1}+B_{2}+B_{3}(E_{0}+K_{0})+B_{4}(D_{0}+K_{0})$.
Proof.
Using (1) and (2) of Lemma 3.1, we have that
$\left\lVert x_{n+2}-x_{n+1}\right\rVert\leq(1-\alpha_{n+1})\left\lVert x_{n+1}-x_{n}\right\rVert+c_{n}$
for all $n\in\mathbb{N}$ everywhere on $\Omega$ where
$c_{n}:=$ $\left\lVert\xi_{n+1}-\xi_{n}\right\rVert+|\alpha_{n+1}-\alpha_{n}|(\left\lVert Tx_{n}-u\right\rVert+\left\lVert\xi_{n}\right\rVert)$
$+\left\lVert\delta_{n+1}-\delta_{n}\right\rVert+|\beta_{n+1}-\beta_{n}|(\left\lVert Uy_{n}-y_{n}\right\rVert+\left\lVert\delta_{n}\right\rVert).$
It is immediate that
$\mathbb{E}[c_{n}]\leq\mathbb{E}[\left\lVert\xi_{n+1}-\xi_{n}\right\rVert]+\mathbb{E}[\left\lVert\delta_{n+1}-\delta_{n}\right\rVert]+|\alpha_{n+1}-\alpha_{n}|(K_{0}+E_{0})+|\beta_{n+1}-\beta_{n}|(K_{0}+D_{0})$
and so
$\chi(\varepsilon):=\max\{\chi_{1}(\varepsilon/4),\chi_{2}(\varepsilon/4),\chi_{3}(\varepsilon/4(E_{0}+K_{0})),\chi_{4}(\varepsilon/4(D_{0}+K_{0}))\}$
is a rate of convergence for $\sum_{n=0}^{\infty}\mathbb{E}[c_{n}]<\infty$, while $B$ as defined above is an upper bound for $\sum_{n=0}^{\infty}\mathbb{E}[c_{n}]$. Naturally, the above yields
$\mathbb{E}[\left\lVert x_{n+1}-x_{n}\right\rVert]\leq\mathbb{E}[\left\lVert x_{1}-x_{0}\right\rVert]+\sum_{i=0}^{n-1}\mathbb{E}[c_{i}]\leq\mathbb{E}[\left\lVert x_{1}-x_{0}\right\rVert]+B$
and we can then show that
$\mathbb{E}[\left\lVert x_{1}-x_{0}\right\rVert]\leq\mathbb{E}[\left\lVert y_{0}-u\right\rVert]+\mathbb{E}[\left\lVert Uu-u\right\rVert]+\mathbb{E}[\left\lVert\xi_{0}\right\rVert]+\mathbb{E}[\left\lVert\delta_{0}\right\rVert]\leq 2K_{0}+E_{0}+D_{0}$
so that $\mathbb{E}[\left\lVert x_{n+1}-x_{n}\right\rVert]\leq K:=2K_{0}+E_{0}+D_{0}+B$. Using Lemma 2.4, we then get the desired rates. ∎
We can then immediately transfer that rate to the complementary sequence $(y_{n})$:
Theorem 3.3.
Under the assumptions of Theorem 3.2, we have that $\mathbb{E}[\left\lVert y_{n+1}-y_{n}\right\rVert]\to 0$ with rate
$\varphi^{\prime}(\varepsilon):=\max\{\varphi(\varepsilon/3),\chi_{1}(\varepsilon/3),\chi_{3}(\varepsilon/3(K_{0}+E_{0}))\}.$
as well as $\left\lVert y_{n+1}-y_{n}\right\rVert\to 0$ almost surely with rate
$\psi^{\prime}(\lambda,\varepsilon):=\max\left\{\varphi^{\prime}(\lambda\varepsilon/2),\chi(\lambda\varepsilon/2)\right\}$
where $\varphi$ is a rate for $\mathbb{E}[\left\lVert x_{n+1}-x_{n}\right\rVert]\to 0$ and $\chi$ is as in Theorem 3.2.
Proof.
Using (1) of Lemma 3.1, we get that
$\left\lVert y_{n+1}-y_{n}\right\rVert\leq\left\lVert x_{n+1}-x_{n}\right\rVert+\left\lVert\xi_{n+1}-\xi_{n}\right\rVert+|\alpha_{n+1}-\alpha_{n}|(\left\lVert Tx_{n}-u\right\rVert+\left\lVert\xi_{n}\right\rVert)$
everywhere on $\Omega$ and for any $n\in\mathbb{N}$. Under expectation, we thus have
$\mathbb{E}[\left\lVert y_{n+1}-y_{n}\right\rVert]\leq\mathbb{E}[\left\lVert x_{n+1}-x_{n}\right\rVert]+\mathbb{E}[\left\lVert\xi_{n+1}-\xi_{n}\right\rVert]+|\alpha_{n+1}-\alpha_{n}|(K_{0}+E_{0})$
and from that the rate for $\mathbb{E}[\left\lVert y_{n+1}-y_{n}\right\rVert]\to 0$ immediately follows, noting that a rate of convergence $\chi$ for a series $\sum_{n=0}^{\infty}a_{n}<\infty$ yields that $\sum_{n=\chi(\varepsilon)}^{\infty}a_{n}<\varepsilon$ and so implies that $a_{n}<\varepsilon$ for any $n\geq\chi(\varepsilon)$. For the rate of $\left\lVert y_{n+1}-y_{n}\right\rVert\to 0$ almost surely, note that using both (1) and (2) of Lemma 3.1, we get that
$\left\lVert y_{n+2}-y_{n+1}\right\rVert\leq\left\lVert y_{n+1}-y_{n}\right\rVert+d_{n}$
for all $n\in\mathbb{N}$ everywhere on $\Omega$ where
$d_{n}:=$ $\left\lVert\xi_{n+2}-\xi_{n+1}\right\rVert+|\alpha_{n+2}-\alpha_{n+1}|(\left\lVert Tx_{n+1}-u\right\rVert+\left\lVert\xi_{n+1}\right\rVert)$
$+\left\lVert\delta_{n+1}-\delta_{n}\right\rVert+|\beta_{n+1}-\beta_{n}|(\left\lVert Uy_{n}-y_{n}\right\rVert+\left\lVert\delta_{n}\right\rVert).$
So, it is immediate that we have
$\mathbb{E}[d_{n}]\leq\mathbb{E}[\left\lVert\xi_{n+2}-\xi_{n+1}\right\rVert]+\mathbb{E}[\left\lVert\delta_{n+1}-\delta_{n}\right\rVert]+|\alpha_{n+2}-\alpha_{n+1}|(K_{0}+E_{0})+|\beta_{n+1}-\beta_{n}|(K_{0}+D_{0}).$
and so that $\chi$ from Theorem 3.2 is a rate of convergence for $\sum_{n=0}^{\infty}\mathbb{E}[d_{n}]<\infty$ (noting that if $\chi$ is a rate of convergence for $\sum_{n=0}^{\infty}a_{n}<\infty$, then $\sum_{n=\chi(\varepsilon)}a_{n+1}=\sum_{n=\chi(\varepsilon)+1}a_{n}\leq\sum_{n=\chi(\varepsilon)}a_{n}<\varepsilon$ so that $\chi$ is also a rate of convergence for $\sum_{n=0}^{\infty}a_{n+1}<\infty$). Using Lemma 2.3, we get the desired rate for $\left\lVert y_{n+1}-y_{n}\right\rVert\to 0$ almost surely. ∎
3.2. Asymptotic regularity relative to the mappings
We now move on to establishing rates of asymptotic regularity for the iterations relative to the mappings. For that, we will actually see a crucial dichotomy, where results based on the use of just one of the mappings $U$ or $T$ are comparatively straightforward, whereas for the general case where neither $U$ nor $T$ trivialize, we rely on a geometric argument for establishing a rate of asymptotic regularity relative to $U$ for the sequence $(y_{n})$, which requires both a uniform convexity of the space $X$ along with a uniform integrability assumption on $(\left\lVert Uy_{n}-y_{n}\right\rVert)$. As such, before we move on to these results, we first give rates of asymptotic regularity relative to the mappings for the remaining cases, dependent on the relevant rates for $(\left\lVert Uy_{n}-y_{n}\right\rVert)$:
Theorem 3.4.
Assume that
$\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$
with rate $\varphi$, and $\mathbb{E}[\left\lVert\xi_{n}\right\rVert],\mathbb{E}[\left\lVert\delta_{n}\right\rVert],\alpha_{n}\to 0$ with rates $\rho_{1}$$\rho_{3}$, respectively. Assume further that $\mathbb{E}[\left\lVert x_{n+1}-x_{n}\right\rVert]\to 0$ with a rate $\varphi_{0}$. Then
  1. (a)
    $\mathbb{E}[\left\lVert x_{n}-y_{n}\right\rVert]\to 0$ with rate
    $\varphi_{1}(\varepsilon):=\max\left\{\varphi_{0}(\varepsilon/3),\varphi(\varepsilon/3),\rho_{2}(\varepsilon/3)\right\},$
  2. (b)
    $\mathbb{E}[\left\lVert Ty_{n}-y_{n}\right\rVert]\to 0$ with rate
    $\varphi_{2}(\varepsilon):=\max\left\{\varphi_{1}(\varepsilon/3),\rho_{3}(\varepsilon/3K_{0}),\rho_{1}(\varepsilon/3)\right\},$
  3. (c)
    $\mathbb{E}[\left\lVert Ux_{n}-x_{n}\right\rVert]\to 0$ with rate
    $\varphi_{3}(\varepsilon):=\max\left\{\varphi_{1}(\varepsilon/3),\varphi(\varepsilon/3)\right\},$
  4. (d)
    $\mathbb{E}[\left\lVert Tx_{n}-x_{n}\right\rVert]\to 0$ with rate
    $\varphi_{4}(\varepsilon):=\max\left\{\varphi_{1}(\varepsilon/3),\varphi_{2}(\varepsilon/3)\right\}.$
Proof.
Related to (a) – (d), we can immediately establish the following inequalities:
$\left\lVert x_{n}-y_{n}\right\rVert$ $\leq\left\lVert x_{n+1}-x_{n}\right\rVert+\left\lVert x_{n+1}-y_{n}\right\rVert$
$\leq\left\lVert x_{n+1}-x_{n}\right\rVert+\left\lVert Uy_{n}-y_{n}\right\rVert+\left\lVert\delta_{n}\right\rVert,$
$\left\lVert Ty_{n}-y_{n}\right\rVert$ $\leq\left\lVert Ty_{n}-Tx_{n}\right\rVert+\left\lVert Tx_{n}-y_{n}\right\rVert$
$\leq\left\lVert y_{n}-x_{n}\right\rVert+\alpha_{n}\left\lVert Tx_{n}-u\right\rVert+\left\lVert\xi_{n}\right\rVert,$
$\left\lVert Ux_{n}-x_{n}\right\rVert$ $\leq\left\lVert Ux_{n}-Uy_{n}\right\rVert+\left\lVert Uy_{n}-y_{n}\right\rVert+\left\lVert y_{n}-x_{n}\right\rVert$
$\leq 2\left\lVert x_{n}-y_{n}\right\rVert+\left\lVert Uy_{n}-y_{n}\right\rVert,$
$\left\lVert Tx_{n}-x_{n}\right\rVert$ $\leq\left\lVert Tx_{n}-Ty_{n}\right\rVert+\left\lVert Ty_{n}-y_{n}\right\rVert+\left\lVert y_{n}-x_{n}\right\rVert$
$\leq 2\left\lVert x_{n}-y_{n}\right\rVert+\left\lVert Ty_{n}-y_{n}\right\rVert.$
By taking the expectation, the rates immediately follow. ∎
We then can similar give rates of asymptotic regularity almost surely under a slight extension of the previous conditions on the errors:
Theorem 3.5.
Under the assumption (Hyp), assume that
$\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$
almost surely with rate $\psi$, and $\alpha_{n}\to 0$ with rate $\rho$. Further, assume that $\left\lVert\xi_{n}\right\rVert,\left\lVert\delta_{n}\right\rVert\to 0$ almost surely with rates $\phi_{1},\phi_{2}$, respectively, and that $\left\lVert x_{n+1}-x_{n}\right\rVert\to 0$ almost surely with a rate $\psi_{0}$. Then
  1. (a)
    $\left\lVert x_{n}-y_{n}\right\rVert\to 0$ almost surely with rate
    $\psi_{1}(\lambda,\varepsilon):=\max\left\{\psi_{0}(\lambda/3,\varepsilon/3),\psi(\lambda/3,\varepsilon/3),\phi_{2}(\lambda/3,\varepsilon/3)\right\},$
  2. (b)
    $\left\lVert Ty_{n}-y_{n}\right\rVert\to 0$ almost surely with rate
    $\psi_{2}(\lambda,\varepsilon):=\max\left\{\psi_{1}(\lambda/3,\varepsilon/3),\rho(\varepsilon\lambda/9K_{0}),\phi_{1}(\lambda/3,\varepsilon/3)\right\},$
  3. (c)
    $\left\lVert Ux_{n}-x_{n}\right\rVert\to 0$ almost surely with rate
    $\psi_{3}(\lambda,\varepsilon):=\max\left\{\psi_{1}(\lambda/2,\varepsilon/3),\psi(\lambda/2,\varepsilon/3)\right\},$
  4. (d)
    $\left\lVert Tx_{n}-x_{n}\right\rVert\to 0$ almost surely with rate
    $\psi_{4}(\lambda,\varepsilon):=\max\left\{\psi_{1}(\lambda/2,\varepsilon/3),\psi_{2}(\lambda/2,\varepsilon/3)\right\}.$
Proof.
The results follow immediately by the same inequalities established in the proof of Theorem 3.4 where in the case (b) one just needs the following additional consideration, giving a rate for $\alpha_{n}\left\lVert Tx_{n}-u\right\rVert\to 0$ almost surely: Using Markov’s inequality we have
$\mathbb{P}(\exists n\left(\left\lVert Tx_{n}-u\right\rVert\geq K_{0}/\lambda\right))\leq\mathbb{P}(Y\geq K_{0}/\lambda)\leq\lambda$
for $Y$ as in (Hyp). Now noting that if $\omega$ is such that $\left\lVert Tx_{n}(\omega)-u(\omega)\right\rVert<K_{0}/\lambda$ for all $n\in\mathbb{N}$, then $\alpha_{n}\left\lVert Tx_{n}(\omega)-u(\omega)\right\rVert<\varepsilon$ for any $n\geq\rho(\varepsilon\lambda/K_{0})$, and therefore we have
$\mathbb{P}(\exists n\geq\rho(\varepsilon\lambda/K_{0})(\alpha_{n}\left\lVert Tx_{n}-u\right\rVert\geq\varepsilon))\leq\mathbb{P}(\exists n\left(\left\lVert Tx_{n}-u\right\rVert\geq K_{0}/\lambda\right))\leq\lambda$
for any $\varepsilon,\lambda>0$ and this suffices to establish the claim in this case. ∎
Remark 3.6.
If we assume that
$\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert],\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<\infty$
with rates of convergence $\chi_{1}$, $\chi_{2}$, respectively, then we can immediately derive rates $\phi_{1},\phi_{2}$ for $\left\lVert\xi_{n}\right\rVert,\left\lVert\delta_{n}\right\rVert\to 0$ almost surely: Since $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]<\infty$ with rate $\chi_{1}$, we get
$\mathbb{P}(\exists n\geq\chi_{1}(\lambda\varepsilon)(\left\lVert\xi_{n}\right\rVert\geq\varepsilon))\leq\sum_{n=\chi_{1}(\lambda\varepsilon)}^{\infty}\mathbb{P}(\left\lVert\xi_{n}\right\rVert\geq\varepsilon)\leq\sum_{n=\chi_{1}(\lambda\varepsilon)}^{\infty}\frac{\mathbb{E}[\left\lVert\xi_{n}\right\rVert]}{\varepsilon}<\lambda$
using Markov’s inequality so that $\phi_{1}(\lambda,\varepsilon):=\chi_{1}(\lambda\varepsilon)$ is a rate for $\left\lVert\xi_{n}\right\rVert\to 0$ almost surely. Similarly for $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<\infty$ and $\chi_{2},\phi_{2}$.
3.2.1. Special cases of the Halpern iteration and the Krasnoselskii-Mann iteration with Tikhonov regularization terms
In the special case of $U:=\mathrm{Id}$ and $\delta_{n}:=0$, the iteration (1) collapses to the stochastic Halpern iteration (1). We then have trivial rates for $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ and $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely and so, in that case, we get under the assumptions of Theorems 3.2 and 3.4 (and also under suitable monotonicity assumptions of the rates involved) that
$\mathbb{E}[\left\lVert Tx_{n}-x_{n}\right\rVert]\to 0$
with a rate
$\varphi(\varepsilon):=\max\{\varphi_{K,\theta,\chi}(\varepsilon/27),\rho_{3}(\varepsilon/9K_{0}),\rho_{1}(\varepsilon/9)\}$
with $\varphi_{K,\theta,\chi}$ defined as in Theorem 3.2 and $\rho_{1},\rho_{3}$ as in Theorem 3.4. Note that this generalises known rates in this case [4, Theorem 3.3], which apply only to specific choices of the parameters. In a similar way, we get a new rate for $\left\lVert Tx_{n}-x_{n}\right\rVert\to 0$ almost surely, though we do not spell it out here.
In the special case of $T:=\mathrm{Id}$ and $\xi_{n}:=0$, the iteration (1) collapses to a stochastic variant of the Krasnoselskii-Mann iteration with Tikhonov regularization terms (1) (and even a slight extension by allowing general anchors $u$). In that case, we do not need to rely on the geometric arguments discussed in the next part of this section and can instead directly derive rates of convergence for $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ and $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely, essentially following the approach of [38] (see also [6]).
Lemma 3.7 (essentially [38]).
Let $(x_{n}),(y_{n})$ be the sequences generated by (1) where $T=\mathrm{Id}$ and $\xi_{n}=0$ for all $n\in\mathbb{N}$. Then the following recurrence relation holds everywhere on $\Omega$ for all $n\in\mathbb{N}$:
$\left\lVert Uy_{n+1}-y_{n+1}\right\rVert\leq 2\left\lVert y_{n}-y_{n+1}\right\rVert+\alpha_{n+1}\left\lVert Uy_{n+1}-u\right\rVert+\left\lVert\delta_{n}\right\rVert+\beta_{n}\left\lVert Uy_{n+1}-y_{n+1}\right\rVert.$
Proof.
Given $n\in\mathbb{N}$, we have
$\left\lVert Uy_{n+1}-y_{n+1}\right\rVert$ $=\left\lVert Uy_{n+1}-(1-\alpha_{n+1})x_{n+1}-\alpha_{n+1}u\right\rVert$
$\leq\left\lVert Uy_{n+1}-x_{n+1}\right\rVert+\alpha_{n+1}\left\lVert Uy_{n+1}-u\right\rVert$
$\leq\left\lVert Uy_{n+1}-Uy_{n}\right\rVert+\beta_{n}\left\lVert Uy_{n+1}-y_{n}\right\rVert+\left\lVert\delta_{n}\right\rVert+\alpha_{n+1}\left\lVert Uy_{n+1}-u\right\rVert$
$\leq 2\left\lVert y_{n+1}-y_{n}\right\rVert+\beta_{n}\left\lVert Uy_{n+1}-y_{n+1}\right\rVert+\left\lVert\delta_{n}\right\rVert+\alpha_{n+1}\left\lVert Uy_{n+1}-u\right\rVert$
pointwise everywhere. ∎
From that inequality, the following rates follow in a straightforward way:
Theorem 3.8.
Let $(x_{n}),(y_{n})$ be the sequences generated by (1) where $T=\mathrm{Id}$ and $\xi_{n}=0$ for all $n\in\mathbb{N}$. Also, let $\Lambda>0$ be such that $\Lambda\leq\beta_{n}\leq 1-\Lambda$ for all $n\in\mathbb{N}$. If $\mathbb{E}[\left\lVert y_{n}-y_{n+1}\right\rVert]\to 0$ with rate $\varphi$, $\alpha_{n}\to 0$ with rate $\rho$ and $\mathbb{E}[\left\lVert\delta_{n}\right\rVert]\to 0$ with rate $\chi$, then $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ with rate
$\kappa(\varepsilon):=\max\{\varphi(\Lambda\varepsilon/4),\rho(\Lambda\varepsilon/4K_{0}),\chi(\Lambda\varepsilon/4)\}+1.$
Under the alternative hypothesis (Hyp) and assuming $\left\lVert y_{n}-y_{n+1}\right\rVert\to 0$ almost surely with rate $\psi$, $\alpha_{n}\to 0$ with rate $\rho$ and $\left\lVert\delta_{n}\right\rVert\to 0$ almost surely with rate $\phi$, then $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely with rate
$\zeta(\lambda,\varepsilon):=\max\{\psi(\lambda/3,\Lambda\varepsilon/4),\rho(\Lambda\lambda\varepsilon/4K_{0}),\phi(\lambda/3,\Lambda\varepsilon/4)\}+1.$
Proof.
By the Lemma 3.7 above, we have
$(1-\beta_{n})\left\lVert Uy_{n+1}-y_{n+1}\right\rVert\leq 2\left\lVert y_{n}-y_{n+1}\right\rVert+\alpha_{n+1}\left\lVert Uy_{n+1}-u\right\rVert+\left\lVert\delta_{n}\right\rVert$
pointwise everywhere for any $n\in\mathbb{N}$. After taking expectations, we have
$\mathbb{E}[\left\lVert Uy_{n+1}-y_{n+1}\right\rVert]\leq\frac{1}{\Lambda}\left(2\mathbb{E}[\left\lVert y_{n}-y_{n+1}\right\rVert]+\alpha_{n+1}K_{0}+\mathbb{E}[\left\lVert\delta_{n}\right\rVert]\right)$
and from that the first rate immediately follows. The second rate follows rather similarly from the above relation: Using Markov’s inequality, it holds that
$\mathbb{P}\left(\exists n\left(\left\lVert Uy_{n+1}-u\right\rVert\geq\frac{K_{0}}{\lambda}\right)\right)\leq\mathbb{P}\left(Y\geq\frac{K_{0}}{\lambda}\right)\leq\lambda$
for all $\lambda>0$. Let now $\lambda,\varepsilon>0$ be given. Take then $\omega$ such that $\left\lVert y_{n}(\omega)-y_{n+1}(\omega)\right\rVert<\Lambda\varepsilon/4$ for all $n\geq\psi(\lambda/3,\Lambda\varepsilon/4)$ and $\left\lVert\delta_{n}(\omega)\right\rVert<\Lambda\varepsilon/4$ for all $n\geq\phi(\lambda/3,\Lambda\varepsilon/4)$ as well as $\left\lVert Uy_{n+1}(\omega)-u(\omega)\right\rVert\leq\frac{K_{0}}{\lambda}$ for all $n$. Then for $n\geq\zeta(\lambda,\varepsilon)-1$, it follows from the above inequality that
$\left\lVert Uy_{n+1}(\omega)-y_{n+1}(\omega)\right\rVert<\varepsilon.$
Therefore we immediately have
$\mathbb{P}\left(\exists n\geq\zeta(\lambda,\varepsilon)\left(\left\lVert Uy_{n}-y_{n}\right\rVert\geq\varepsilon\right)\right)<\lambda$
which completes the proof. ∎
3.2.2. The general case
[…]
$D\geq\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]\text{ and }E\geq\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$
as well as a $K_{0}$ such that $K_{0}\geq\mathbb{E}[\left\lVert x_{0}-p\right\rVert],\mathbb{E}[\left\lVert u-p\right\rVert]$. Using these data, we immediately get the following extended result on bounds:
Lemma 3.9.
For all $n\in\mathbb{N}$, $\left\lVert x_{n}-p\right\rVert\leq Y^{\prime}\leq 2Y^{\prime}=:Y$ pointwise everywhere, where
$Y^{\prime}:=\left\lVert x_{0}-p\right\rVert+\left\lVert u-p\right\rVert+\sum_{i=0}^{\infty}\left(\left\lVert\xi_{n}\right\rVert+\left\lVert\delta_{n}\right\rVert\right)$
and furthermore $\mathbb{E}[Y]\leq K:=4K_{0}+2D+2E$. The sequences
$\left\lVert y_{n}-p\right\rVert,\left\lVert x_{n+1}-x_{n}\right\rVert,\left\lVert y_{n+1}-y_{n}\right\rVert,\left\lVert Tx_{n}-u\right\rVert,\left\lVert Uy_{n}-y_{n}\right\rVert,\left\lVert Uy_{n}-u\right\rVert$
are “$L_{1}$-dominated” by $Y$ in a similar way.
Proof.
Pointwise everywhere it holds that
$\left\lVert x_{n+1}-p\right\rVert$ $=\left\lVert(1-\beta_{n})(Uy_{n}+\delta_{n})+\beta_{n}y_{n}-p\right\rVert$
$\leq(1-\beta_{n})\left\lVert Uy_{n}-p\right\rVert+\beta_{n}\left\lVert y_{n}-p\right\rVert+\left\lVert\delta_{n}\right\rVert$
$\leq\left\lVert y_{n}-p\right\rVert+\left\lVert\delta_{n}\right\rVert$
$=\left\lVert(1-\alpha_{n})(Tx_{n}+\xi_{n})+\alpha_{n}u-p\right\rVert+\left\lVert\delta_{n}\right\rVert$
$\leq(1-\alpha_{n})\left\lVert Tx_{n}-p\right\rVert+\alpha_{n}\left\lVert u-p\right\rVert+\left\lVert\xi_{n}\right\rVert+\left\lVert\delta_{n}\right\rVert$
$\leq(1-\alpha_{n})\left\lVert x_{n}-p\right\rVert+\alpha_{n}\left\lVert u-p\right\rVert+\left\lVert\xi_{n}\right\rVert+\left\lVert\delta_{n}\right\rVert.$
It follows immediately by induction that
$\left\lVert x_{n+1}-p\right\rVert\leq Y^{\prime}_{n}:=\left\lVert x_{0}-p\right\rVert+\left\lVert u-p\right\rVert+\sum_{i=0}^{n}\left(\left\lVert\xi_{n}\right\rVert+\left\lVert\delta_{n}\right\rVert\right)$
holds pointwise everywhere. Since the $Y^{\prime}_{n}$ are pointwise monotone, defining $Y^{\prime}:=\sup_{n\in\mathbb{N}}Y^{\prime}_{n}$ yields that $\left\lVert x_{n}-p\right\rVert\leq Y^{\prime}_{n}\leq Y^{\prime}$ pointwise everywhere for all $n\in\mathbb{N}$, and by the monotone convergence theorem we have
$\mathbb{E}[Y^{\prime}]=\mathbb{E}[\left\lVert x_{0}-p\right\rVert]+\mathbb{E}[\left\lVert u-p\right\rVert]+\sum_{i=0}^{\infty}\left(\mathbb{E}[\left\lVert\xi_{n}\right\rVert]+\mathbb{E}[\left\lVert\delta_{n}\right\rVert]\right)\leq 2K_{0}+D+E.$
Therefore immediately $\mathbb{E}[Y]\leq K$. By the above inequalities, one also has $\left\lVert y_{n}-p\right\rVert\leq Y^{\prime}_{n}\leq Y^{\prime}$, and the rest of the bounds follow by the triangle inequality. ∎
For the rest of this section, we will always assume the existence of a fixed point as above and use $Y$ and $K$ to refer to the quantities in Lemma 3.9. Note that these in particular validate the assumptions (Hyp) and (Hyp).
We now move on to the asymptotic regularity relative to $U$ of the sequence $(y_{n})$. For that, we initially establish $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely using geometric properties of the underlying space and then use quantitative uniform integrability assumption for this random variables to recover the asymptotic regularity relative to $U$ of $(y_{n})$ in expectation. First, we will make the following geometric assumption on the underlying normed space $(X,\left\lVert\cdot\right\rVert)$:
Definition 3.10.
We say that $(X,\left\lVert\cdot\right\rVert)$ is uniformly convex ([9]) if for any $\varepsilon\in(0,2]$, there exists a $\delta\in(0,1]$ such that for all $x,y\in\overline{B}_{1}(0)$:
$\left\lVert x-y\right\rVert\geq\varepsilon\text{ implies }\left\lVert\frac{x+y}{2}\right\rVert\leq 1-\delta.$
We call a modulus $\eta:(0,2]\to(0,1]$ witnessing such a $\delta$ in terms of $\varepsilon$ a modulus of uniform convexity for $X$.
The above modulus also applies to closed balls of any radius centered at any point in the space and for arbitrary convex combinations:
Lemma 3.11.
Let $\eta$ be a modulus of uniform convexity. For any $r>0$ and $\varepsilon\in(0,2]$, if $x,y\in\overline{B}_{r}(a)$ for $a\in X$ with $\left\lVert x-y\right\rVert\geq\varepsilon\cdot r$, then for all $\lambda\in[0,1]$:
$\left\lVert(1-\lambda)x+\lambda y-a\right\rVert\leq(1-2\lambda(1-\lambda)\eta(\varepsilon))r.$
The proof is straightforward and we hence omit it (but refer e.g. [27] for a proof of such a property even in the context of nonlinear uniformly convex hyperbolic spaces).
The proof of the following theorem now follows the outline of the proof of an analogous result for the Halpern-Mann iteration in uniformly convex hyperbolic spaces as given in [29] (though without errors, even nonstochastic ones):
Theorem 3.12.
Let $(X,\left\lVert\cdot\right\rVert)$ be uniformly convex with modulus $\eta$. Let $\left\lVert x_{n+1}-x_{n}\right\rVert\to 0$ almost surely with rate $\Delta$. Also, let $\rho$ be a rate for $\alpha_{n}\to 0$ and assume that
$\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert],\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<\infty$
with rates of convergence $\chi_{1}$, $\chi_{2}$, respectively. Lastly, let $\Lambda>0$ be such that $\Lambda\leq\beta_{n}\leq 1-\Lambda$. Then $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely with rate
$\Gamma(\lambda,\varepsilon):=\max\{\Delta(\lambda/9,\widehat{\varepsilon}/4),\rho(\widehat{\varepsilon}/4K^{\prime}),\chi_{1}(\lambda\widehat{\varepsilon}/36),\chi_{2}(\lambda\widehat{\varepsilon}/36)\}$
for $\widehat{\varepsilon}:=\varepsilon\cdot\Lambda^{2}\cdot\eta(\varepsilon/K^{\prime})$ and $K^{\prime}:=3K/\lambda$.
Proof.
Suppose for contradiction that
$\mathbb{P}\left(\exists n\geq\Gamma(\lambda,\varepsilon)\,(\left\lVert Uy_{n}-y_{n}\right\rVert\geq\varepsilon)\right)\geq\lambda$
and call the set inside the probability $B_{\lambda,\varepsilon}$. By Lemma 3.9 and Markov’s inequality, we have
$\mathbb{P}\left(\exists n\left(\left\lVert y_{n}-p\right\rVert>\frac{K}{\lambda}\right)\right)\leq\mathbb{P}\left(Y\geq\frac{K}{\lambda}\right)\leq\frac{\mathbb{E}[Y]}{K/\lambda}\leq\lambda$
for any $\lambda>0$ and so $\mathbb{P}\left(\exists n\left(\left\lVert y_{n}-p\right\rVert>K^{\prime}\right)\right)\leq\lambda/3$ for $K^{\prime}:=3K/\lambda$. Similarly for $\left\lVert u-p\right\rVert$. Thus, using the Fréchet inequalities, we have
$\mathbb{P}(\exists n\geq\Gamma(\lambda,\varepsilon)(\left\lVert Uy_{n}-y_{n}\right\rVert\geq\varepsilon)\land\forall n\left(\left\lVert y_{n}-p\right\rVert,\left\lVert u-p\right\rVert\leq K^{\prime}\right))$
$\qquad\geq\mathbb{P}(\exists n\geq\Gamma(\lambda,\varepsilon)(\left\lVert Uy_{n}-y_{n}\right\rVert>\varepsilon))+\mathbb{P}(\forall n(\left\lVert y_{n}-p\right\rVert\leq K^{\prime}))+\mathbb{P}(\forall n(\left\lVert u-p\right\rVert\leq K^{\prime}))-2$
$\qquad\geq\lambda+(1-\lambda/3)+(1-\lambda/3)-2$
$\qquad=\lambda/3.$
We denote that set measured in the above by $A_{\lambda,\varepsilon}$, and let $\omega\in A_{\lambda,\varepsilon}$, i.e. there exists some $n(\omega)\geq\Gamma(\lambda,\varepsilon)$ such that
$\left\lVert Uy_{n(\omega)}(\omega)-y_{n(\omega)}(\omega)\right\rVert\geq\varepsilon\text{ and }\left\lVert y_{n(\omega)}(\omega)-p\right\rVert,\left\lVert u(\omega)-p\right\rVert\leq K^{\prime}.$
Writing $n_{0}$ for $n(\omega)$, we then have $\left\lVert Uy_{n_{0}}(\omega)-y_{n_{0}}(\omega)\right\rVert\leq 2\left\lVert y_{n_{0}}(\omega)-p\right\rVert\leq 2K^{\prime}$ so that
$\varepsilon/2\leq\left\lVert y_{n_{0}}(\omega)-p\right\rVert\leq K^{\prime}.$
Also, we have $\left\lVert Uy_{n_{0}}(\omega)-p\right\rVert\leq\left\lVert y_{n_{0}}(\omega)-p\right\rVert\leq K^{\prime}$ as well as
$\left\lVert Uy_{n_{0}}(\omega)-y_{n_{0}}(\omega)\right\rVert\geq\varepsilon=\varepsilon/K^{\prime}\cdot K^{\prime}\geq\varepsilon/K^{\prime}\left\lVert y_{n_{0}}(\omega)-p\right\rVert$
and $\varepsilon/K^{\prime}\leq 2$. So, we can apply Lemma 3.11 to derive
$\left\lVert x_{{n_{0}}+1}(\omega)-p\right\rVert$ $=\left\lVert(1-\beta_{n_{0}})(Uy_{n_{0}}(\omega)+\delta_{n_{0}}(\omega))+\beta_{n_{0}}y_{n_{0}}(\omega)-p\right\rVert$
$\leq\left\lVert(1-\beta_{n_{0}})Uy_{n_{0}}(\omega)+\beta_{n_{0}}y_{n_{0}}(\omega)-p\right\rVert+\left\lVert\delta_{n_{0}}(\omega)\right\rVert$
$\leq(1-2\beta_{n_{0}}(1-\beta_{n_{0}})\eta(\varepsilon/K^{\prime}))\left\lVert y_{n_{0}}(\omega)-p\right\rVert+\left\lVert\delta_{n_{0}}(\omega)\right\rVert$
$\leq\left\lVert y_{n_{0}}(\omega)-p\right\rVert-2\left\lVert y_{n_{0}}(\omega)-p\right\rVert\Lambda^{2}\eta(\varepsilon/K^{\prime})+\left\lVert\delta_{n_{0}}(\omega)\right\rVert$
$\leq\left\lVert y_{n_{0}}(\omega)-p\right\rVert-\varepsilon\cdot\Lambda^{2}\cdot\eta(\varepsilon/K^{\prime})+\left\lVert\delta_{n_{0}}(\omega)\right\rVert.$
Now, we further have
$\left\lVert y_{n_{0}}(\omega)-p\right\rVert$ $\leq(1-\alpha_{n_{0}})\left\lVert Tx_{n_{0}}(\omega)-p\right\rVert+\alpha_{n_{0}}\left\lVert u(\omega)-p\right\rVert+\left\lVert\xi_{n_{0}}(\omega)\right\rVert$
$\leq\left\lVert x_{n_{0}}(\omega)-p\right\rVert+\alpha_{n_{0}}\left\lVert u(\omega)-p\right\rVert+\left\lVert\xi_{n_{0}}(\omega)\right\rVert$
so that we can in particular derive
$\left\lVert x_{{n_{0}}+1}(\omega)-p\right\rVert\leq\left\lVert x_{n_{0}}(\omega)-p\right\rVert+\alpha_{n_{0}}K^{\prime}+\left\lVert\xi_{n_{0}}(\omega)\right\rVert+\left\lVert\delta_{n_{0}}(\omega)\right\rVert-\varepsilon\cdot\Lambda^{2}\cdot\eta(\varepsilon/K^{\prime}).$
So, in the end we have
$\widehat{\varepsilon}$ $=\varepsilon\cdot\Lambda^{2}\cdot\eta(\varepsilon/K^{\prime})$
$\leq\left\lVert x_{n_{0}}(\omega)-p\right\rVert-\left\lVert x_{{n_{0}}+1}(\omega)-p\right\rVert+\alpha_{n_{0}}K^{\prime}+\left\lVert\xi_{n_{0}}(\omega)\right\rVert+\left\lVert\delta_{n_{0}}(\omega)\right\rVert$
$\leq\left\lVert x_{{n_{0}}+1}(\omega)-x_{n_{0}}(\omega)\right\rVert+\alpha_{n_{0}}K^{\prime}+\left\lVert\xi_{n_{0}}(\omega)\right\rVert+\left\lVert\delta_{n_{0}}(\omega)\right\rVert.$
Letting $V_{n}:=\left\lVert x_{n+1}-x_{n}\right\rVert+\alpha_{n}K^{\prime}+\left\lVert\xi_{n}\right\rVert+\left\lVert\delta_{n}\right\rVert$, we have shown that
$A_{\lambda,\varepsilon}\subseteq\{\exists n\geq\Gamma(\lambda,\varepsilon)\,(V_{n}\geq\widehat{\varepsilon})\}.$
Similarly to in the proof of Theorem 3.5, we now have that $\chi_{1}(\lambda\varepsilon),\chi_{2}(\lambda\varepsilon)$ are rates for $\left\lVert\xi_{n}\right\rVert,\left\lVert\delta_{n}\right\rVert\to 0$, respectively. So we have
$\lambda/3$ $\leq\mathbb{P}(A_{\lambda,\varepsilon})$
$\leq\mathbb{P}(\exists n\geq\Gamma(\lambda,\varepsilon)\,(V_{n}\geq\widehat{\varepsilon}))$
$\leq\mathbb{P}\left(\exists n\geq\Gamma(\lambda,\varepsilon)\,\left((\left\lVert x_{n+1}-x_{n}\right\rVert\geq\widehat{\varepsilon}/4)\cup(\alpha_{n}K^{\prime}\geq\widehat{\varepsilon}/4)\cup(\left\lVert\xi_{n}\right\rVert\geq\widehat{\varepsilon}/4)\cup(\left\lVert\delta_{n}\right\rVert\geq\widehat{\varepsilon}/4)\right)\right)$
$\leq\mathbb{P}\left(\exists n\geq\Delta(\lambda/9,\widehat{\varepsilon}/4)(\left\lVert x_{n+1}-x_{n}\right\rVert\geq\widehat{\varepsilon}/4)\right)+\mathbb{P}\left(\exists n\geq\rho(\widehat{\varepsilon}/4K^{\prime})(\alpha_{n}K^{\prime}\geq\widehat{\varepsilon}/4)\right)$
$\quad+\mathbb{P}\left(\exists n\geq\chi_{1}(\lambda\widehat{\varepsilon}/36)(\left\lVert\xi_{n}\right\rVert\geq\widehat{\varepsilon}/4)\right)+\mathbb{P}\left(\exists n\geq\chi_{2}(\lambda\widehat{\varepsilon}/36)(\left\lVert\delta_{n}\right\rVert\geq\widehat{\varepsilon}/4)\right)$
$<\frac{\lambda}{9}+0+\frac{\lambda}{9}+\frac{\lambda}{9}=\frac{\lambda}{3},$
a contradiction. ∎
Remark 3.13.
Using a slightly different argument first devised in [18, Theorem 3.4] (see also [25, Remark 15] or [29, Remark 3.7] for similar remarks in the context of nonlinear spaces), we can slightly optimize the above rate in the context of moduli of uniform convexity of a special form: Let $\eta(\varepsilon)=\varepsilon\cdot\tilde{\eta}(\varepsilon)$ where $\tilde{\eta}$ is increasing. Then above rate $\Gamma$ holds even with $\widehat{\varepsilon}$ defined as
$\widehat{\varepsilon}:=\varepsilon\cdot\Lambda^{2}\cdot\tilde{\eta}(\varepsilon/K^{\prime}).$
To see this, follow the proof of Theorem 3.12 but replace $\varepsilon/K^{\prime}$ with $\varepsilon/\left\lVert y_{n_{0}}(\omega)-p\right\rVert$. Then also $\varepsilon/\left\lVert y_{n_{0}}(\omega)-p\right\rVert\leq 2$ as well as
$\left\lVert Uy_{n_{0}}(\omega)-y_{n_{0}}(\omega)\right\rVert\geq\varepsilon/\left\lVert y_{n_{0}}(\omega)-p\right\rVert\cdot\left\lVert y_{n_{0}}(\omega)-p\right\rVert$
and this leads to
$\left\lVert x_{{n_{0}}+1}(\omega)-p\right\rVert$ $\leq\left\lVert y_{n_{0}}(\omega)-p\right\rVert-2\cdot\varepsilon\cdot\Lambda^{2}\cdot\tilde{\eta}(\varepsilon/\left\lVert y_{n_{0}}(\omega)-p\right\rVert)+\left\lVert\delta_{n_{0}}(\omega)\right\rVert$
$\leq\left\lVert y_{n_{0}}(\omega)-p\right\rVert-\varepsilon\cdot\Lambda^{2}\cdot\tilde{\eta}(\varepsilon/\left\lVert y_{n_{0}}(\omega)-p\right\rVert)+\left\lVert\delta_{n_{0}}(\omega)\right\rVert$
$\leq\left\lVert y_{n_{0}}(\omega)-p\right\rVert-\varepsilon\cdot\Lambda^{2}\cdot\tilde{\eta}(\varepsilon/K^{\prime})+\left\lVert\delta_{n_{0}}(\omega)\right\rVert.$
using that $\tilde{\eta}(\varepsilon/K^{\prime})\leq\tilde{\eta}(\varepsilon/\left\lVert y_{n_{0}}(\omega)-p\right\rVert)$ as $\tilde{\eta}(\varepsilon/\left\lVert y_{n_{0}}(\omega)-p\right\rVert)\leq K^{\prime}$ and since $\tilde{\eta}$ is increasing. Then the proof continuous as before.
We now discuss the assumptions from the quantitative theory of expected values that we require to establish an analogous result on the asymptotic regularity relative to $U$ of $(y_{n})$ in mean.
Definition 3.14.
Let $X$ be an integrable random variable. We call a function $\mu:(0,\infty)\to(0,\infty)$ such that
$\forall\varepsilon>0\,\forall A\in\mathcal{F}\left(\mathbb{P}(A)\leq\mu(\varepsilon)\to\mathbb{E}[|X|1_{A}]\leq\varepsilon\right)$
a modulus of absolute continuity for $X$.
Lemma 3.15.
Let $X$ be an integrable random variable and $\mu$ a modulus of absolute continuity for $X$. For any $a,\varepsilon\in(0,\infty)$, we have that
$\mathbb{E}[|X|]\geq a+\varepsilon\text{ implies }\mathbb{P}(|X|>a)>\mu(\varepsilon/2).$
In particular, we have that $\mathbb{E}[|X|]\geq\varepsilon$ implies $\mathbb{P}(|X|>\varepsilon/2)>\mu(\varepsilon/4)$.
Proof.
Suppose $\mathbb{P}(|X|>a)\leq\mu(\varepsilon/2)$. Then we have
$\mathbb{E}[|X|]=\mathbb{E}[|X|1_{|X|\leq a}]+\mathbb{E}[|X|1_{|X|>a}]\leq a+\varepsilon/2<a+\varepsilon$
which is the claim. ∎
Definition 3.16.
A sequence of random variables $(X_{n})$ is called uniformly integrable if both $\sup_{n\in\mathbb{N}}\mathbb{E}[|X_{n}|]<\infty$ and for any $\varepsilon>0$, there exists a $\delta>0$ such that
$\forall n\in\mathbb{N}\,\forall A\in\mathcal{F}\left(\mathbb{P}(A)\leq\delta\to\mathbb{E}[|X_{n}|1_{A}]\leq\varepsilon\right).$
We call a function $\mu$ that witnesses such a $\delta$ in terms of $\varepsilon$ a modulus of uniform integrability for $(X_{n})$.
Note $\mu$ is a modulus of uniform integrability for $(X_{n})$ exactly when $\mu$ is a modulus of absolute continuity for any $X_{n}$. The main use that a modulus of uniform integrability has for a stochastic process is that with it, we can transfer a rate of almost-sure convergence to a rate of convergence in mean:
Lemma 3.17.
Let $(X_{n})$ be a sequence of nonnegative random variables such that $X_{n}\to 0$ almost surely with rate $\varphi$ and such that $\mu$ is modulus of uniform integrability for $(X_{n})$. Then $\mathbb{E}[X_{n}]\to 0$ with rate
$\Gamma(\varepsilon):=\varphi\left(\mu\left(\frac{\varepsilon}{4}\right),\frac{\varepsilon}{2}\right).$
Proof.
Suppose for contradiction that there exists some $n_{0}\geq\Gamma(\varepsilon)$ with $\mathbb{E}[X_{n_{0}}]\geq\varepsilon$. Then by Lemma 3.15 we have $\mathbb{P}(X_{n_{0}}>\varepsilon/2)>\mu(\varepsilon/4)$ and hence $\mathbb{P}\left(\exists n\geq\Gamma(\varepsilon)(X_{n}\geq\varepsilon/2)\right)\geq\mathbb{P}(X_{n_{0}}>\varepsilon/2)>\mu(\varepsilon/4)$, a contradiction. ∎
Theorem 3.18.
Let $(X,\left\lVert\cdot\right\rVert)$ be uniformly convex with modulus $\eta$. Under the assumptions of Theorem 3.2, let $\mathbb{E}[\left\lVert x_{n+1}-x_{n}\right\rVert]\to 0$ with rate $\Delta$ from Theorem 3.2. Also, let $\rho$ be a rate for $\alpha_{n}\to 0$ and assume that
$\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert],\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<\infty$
with rates of convergence $\chi_{1}$, $\chi_{2}$, respectively. Also, let $\Lambda>0$ be such that $\Lambda\leq\beta_{n}\leq 1-\Lambda$. Lastly, let $\mu$ be a modulus of uniform integrability for $(\left\lVert Uy_{n}-y_{n}\right\rVert)$. Then $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ with rate
$\Gamma(\varepsilon):=\max\{\Delta(\overline{\varepsilon}),\rho(\widehat{\varepsilon}/4K^{\prime}),\chi_{1}(\overline{\varepsilon}),\chi_{2}(\overline{\varepsilon})\}$
where $\overline{\varepsilon}:=\widehat{\varepsilon}\mu(\varepsilon/4)/36$ for $\widehat{\varepsilon}:=\varepsilon/2\cdot\Lambda^{2}\cdot\eta(\varepsilon/2K^{\prime})$ and $K^{\prime}:=3K/\mu(\varepsilon/4)$.
Proof.
By Theorem 3.2, $\Delta(\lambda\varepsilon/2)$ is a rate of almost sure convergence for $\left\lVert x_{n+1}-x_{n}\right\rVert\to 0$. Therefore by Theorem 3.12, a rate of almost sure convergence for $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ is given by
$\tilde{\Gamma}(\lambda,\varepsilon)$ $:=\max\{\Delta(\lambda\tilde{\varepsilon}/36),\rho(\tilde{\varepsilon}/4K^{\prime}),\chi_{1}(\lambda\tilde{\varepsilon}/36),\chi_{2}(\lambda\tilde{\varepsilon}/36)\}$
for $\tilde{\varepsilon}:=\varepsilon\cdot\Lambda^{2}\cdot\eta(\varepsilon/K^{\prime})$ and $K^{\prime}:=3K/\lambda$. By Lemma 3.17 we have $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ with rate
$\Gamma(\varepsilon):=\tilde{\Gamma}\left(\mu\left(\frac{\varepsilon}{4}\right),\frac{\varepsilon}{2}\right)=\max\{\Delta(\overline{\varepsilon}),\rho(\widehat{\varepsilon}/4K^{\prime}),\chi_{1}(\overline{\varepsilon}),\chi_{2}(\overline{\varepsilon})\}$
where $\overline{\varepsilon}:=\widehat{\varepsilon}\mu(\varepsilon/4)/36$ now for $\widehat{\varepsilon}:=\varepsilon/2\cdot\Lambda^{2}\cdot\eta(\varepsilon/2K^{\prime})$ and $K^{\prime}:=3K/\mu(\varepsilon/4)$. ∎
Remark 3.19.
Using Remark 3.13, it follows that also here, if $\eta(\varepsilon)=\varepsilon\cdot\tilde{\eta}(\varepsilon)$ where $\tilde{\eta}$ is increasing, then above rate $\Gamma$ holds even with $\widehat{\varepsilon}$ defined as $\widehat{\varepsilon}:=\varepsilon/2\cdot\Lambda^{2}\cdot\tilde{\eta}(\varepsilon/2K^{\prime})$.
Before moving on to the case of fast rates for special choices of scalars and errors, we briefly discuss some natural conditions under which the previous additional assumption of a modulus of uniform integrability for the sequence $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ can be obtained (cf. [14, Chapter 5.4]), and indicate how a modulus arises from these.
A first common assumption to establish uniform integrability is to assume a uniform finite $p$-th moment for $p>1$. Now, in terms of an upper bound $K$ witnessing that assumption, i.e. $\sup_{n\in\mathbb{N}}\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert^{p}]<K$ for $p>1$, we can actually compute a corresponding modulus $\mu$. To see this, note that for any $\varepsilon>0$, setting $X_{n}:=\left\lVert Uy_{n}-y_{n}\right\rVert$, we have
$\mathbb{E}[|X_{n}|1_{A}]\leq\mathbb{E}[|X_{n}|1_{A\cap(|X_{n}|\leq a)}]+\mathbb{E}[|X_{n}|1_{A\cap(|X_{n}|>a)}]\leq a\mathbb{P}(A)+a^{1-p}\mathbb{E}[|X_{n}|^{p}]\leq a\mathbb{P}(A)+\varepsilon/2$
for $a:=(2K/\varepsilon)^{1/(p-1)}$ and $A\in\mathcal{F}$ and thus a modulus of uniform integrability is given by
$\mu(\varepsilon):=\frac{\varepsilon}{2}\left(\frac{\varepsilon}{2K}\right)^{1/(p-1)}.$
Another growth condition commonly imposed is to assume uniform finite expectation of the random variables under a supercoercive function. Then, in terms of an upper bound $K$ witnessing that property, i.e. $\sup_{n\in\mathbb{N}}\mathbb{E}[g(\left\lVert Uy_{n}-y_{n}\right\rVert)]<K$ where $g:[0,\infty)\to[0,\infty)$ is supercoercive, i.e.
$\frac{g(x)}{x}\to\infty\text{ as }x\to\infty,$
with a rate of divergence $\kappa:(0,\infty)\to(0,\infty)$, i.e.
$\forall a>0\,\forall x\geq\kappa(a)\left(\frac{g(x)}{x}\geq a\right),$
a corresponding modulus $\mu$ can again be easily computed: Similarly to above we have
$\mathbb{E}[|X_{n}|1_{A}]\leq\kappa(a)\mathbb{P}(A)+\mathbb{E}[g\left(|X_{n}|\right)]/a\leq\kappa(a)\mathbb{P}(A)+\varepsilon/2$
for $a:=2K/\varepsilon$ and thus a modulus of uniform integrability is now given by
$\mu(\varepsilon):=\frac{\varepsilon}{2\kappa\left(2K/\varepsilon\right)}.$
Most importantly, we want to show that the required modulus of uniform integrability can be obtained through quantitative integrability assumptions on the errors. Concretely, note that we have $\left\lVert Uy_{n}-y_{n}\right\rVert\leq Y$ pointwise everywhere by Lemma 3.9, so a modulus of integrability for $Y$ immediately yields a modulus of uniform integrability for $(\left\lVert Uy_{n}-y_{n}\right\rVert)$. In particular, if $u$ and $x_{0}$ are chosen to be constant and $K>0$ is such that $\left\lVert x_{0}-p\right\rVert,\left\lVert u-p\right\rVert<K$, and in addition both $\sum_{i=0}^{\infty}\left\lVert\xi_{i}\right\rVert$ and $\sum_{i=0}^{\infty}\left\lVert\delta_{i}\right\rVert$ are integrable with moduli of absolute continuity $\mu_{1}$ and $\mu_{2}$ respectively, by Lemma 3.9 we have
$\mathbb{E}[|X_{n}|1_{A}]\leq 2K\mathbb{P}(A)+2\left(\sum_{i=0}^{\infty}\left(\mathbb{E}[\left\lVert\xi_{i}\right\rVert 1_{A}]+\mathbb{E}[\left\lVert\delta_{i}\right\rVert 1_{A}]\right)\right)$
and thus a modulus of uniform integrability is now given by
$\mu(\varepsilon):=\min\left\{\frac{\varepsilon}{4K},\mu_{1}\left(\frac{\varepsilon}{8}\right),\mu_{2}\left(\frac{\varepsilon}{8}\right)\right\}.$
In this way, we demonstrate that it is possible to shift the integrability condition on $(\left\lVert Uy_{n}-y_{n}\right\rVert)$ to appropriate integrability conditions on the sums of the error terms.
4. Fast rates of asymptotic regularity
In this section, we focus on particular instantiations of the parameters together with suitable growth conditions on the errors that allow for fast rates of asymptotic regularity for the above iteration(s). For that, we begin with some general results on deriving linear rates of convergence for sequences of real numbers satisfying a general recursive inequality and we subsequently extend this to sequences of random variables and utilize these general results then to in turn derive the fast rates. Throughout the section, we will be very explicit about the exact kind of assumptions (i.e. (Hyp), or (Hyp), or the existence of common fixed points in the last part of the previous section ) that are placed on the iterations in question.
4.1. General results on linear rates
We begin with the crucial result on deriving fast rates of convergence for Halpern-style iterations in nonlinear optimization. This result is closely modelled after a seminal lemma by Sabach and Shtern [36], first utilised in the context of proof mining in [6]. Here we formulate the idea behind the lemma in a slightly different style to fit the iterations considered in this paper, and in this way our presentation is closer to the explicit closed-form bounds in [4].
Lemma 4.1 (essentially [36]).
Suppose that $(s_{n}),(c_{n})$ are sequences of nonnegative real numbers satisfying
$s_{n+1}\leq(1-a_{n})s_{n}+c_{n}$
for all $n\in\mathbb{N}$ where $(a_{n})\subseteq[0,1]$. Then for all $m,K\in\mathbb{N}$ we have
$s_{K+m+1}\leq A_{K}^{K+m}s_{K}+\sum_{i=K}^{K+m}A_{i+1}^{K+m}c_{i}$
for $A_{j}^{k}:=\prod_{i=j}^{k}(1-a_{i})$, with $A_{j}^{k}:=1$ for $j>k$. In the special case that $a_{n}:=\alpha_{n+1}$ and $c_{n}\leq(\alpha_{n}-\alpha_{n+1})L$ for some $(\alpha_{n})\subseteq[0,1]$ and $L>0$ we have
$s_{n}\leq\tilde{A}_{1}^{n}s_{0}+L\sum_{i=1}^{n}(\alpha_{i-1}-\alpha_{i})\tilde{A}_{i+1}^{n}$
for all $n\in\mathbb{N}$ where $\tilde{A}_{j}^{k}:=\prod_{i=j}^{k}(1-\alpha_{i})$, with $\tilde{A}_{j}^{k}:=1$ for $j>k$. If we furthermore define $\alpha_{n}:=2/(n+2)$ and assume that $s_{0}\leq L$, then
$s_{n}\leq\frac{2L}{n+2}$
for all $n\in\mathbb{N}$.
Proof.
The first inequality follows for all $m,K\in\mathbb{N}$ as in Lemma 2.2 immediately by induction. For the second part of the above lemma, we note that $A_{j}^{k}=\tilde{A}_{j+1}^{k+1}$ and thus from this first inequality (after setting $K=0$), we have
$s_{n+1}\leq A_{0}^{n}s_{0}+\sum_{i=0}^{n}A_{i+1}^{n}c_{i}=\tilde{A}_{1}^{n+1}s_{0}+\sum_{i=0}^{n}\tilde{A}_{i+2}^{n+1}(\alpha_{i}-\alpha_{i+1})L\leq\tilde{A}_{1}^{n+1}s_{0}+L\sum_{i=1}^{n+1}\tilde{A}_{i+1}^{n+1}(\alpha_{i-1}-\alpha_{i})$
and also $s_{0}=\tilde{A}_{1}^{0}s_{0}$ by definition. For the final part, we observe that
$\tilde{A}_{i}^{n}=\prod_{j=i}^{n}(1-\alpha_{j})=\prod_{j=i}^{n}\frac{j}{j+2}=\frac{i}{i+2}\cdot\frac{i+1}{i+3}\cdot\ldots\cdot\frac{n-1}{n+1}\cdot\frac{n}{n+2}=\frac{i(i+1)}{(n+1)(n+2)}$
for $i\leq n$ (noting that this also holds for $i=n$), and therefore
$s_{n}$ $\leq\frac{2s_{0}}{(n+1)(n+2)}+L\sum_{i=1}^{n}\left(\frac{2}{i+1}-\frac{2}{i+2}\right)\frac{(i+1)(i+2)}{(n+1)(n+2)}$
$=\frac{2}{(n+1)(n+2)}\left(s_{0}+L\sum_{i=1}^{n}\left(\frac{1}{i+1}-\frac{1}{i+2}\right)(i+1)(i+2)\right)$
$\leq\frac{2L}{(n+1)(n+2)}\left(1+\sum_{i=1}^{n}1\right)=\frac{2L}{n+2}$
which completes the proof. ∎
The following is an adaptation of the special case of the previous lemma concerning fast rates to sequences of random variables and as we will see in the following, it assumes a similarly important role for deriving linear rates of almost sure convergence.
Lemma 4.2.
Suppose that $(X_{n})$, $(C_{n})$ are nonnegative stochastic processes satisfying
$X_{n+1}\leq(1-\alpha_{n+1})X_{n}+C_{n}$
almost surely for any $n\in\mathbb{N}$ where $\alpha_{n}:=2/(n+2)$ and where $\mathbb{E}[C_{n}]\leq(\alpha_{n}-\alpha_{n+1})L$ almost surely for all $n\in\mathbb{N}$ where $L\geq\mathbb{E}[X_{0}]$. Then
$\mathbb{E}[X_{n}]\leq\frac{2L}{n+2}\text{ and }\mathbb{P}\left(\exists i\geq n\left(X_{i}\geq\varepsilon\right)\right)\leq\frac{1}{\varepsilon}\frac{4L}{n+2}$
for all $n\in\mathbb{N}$.
Proof.
From the fact that $X_{n+1}\leq(1-\alpha_{n+1})X_{n}+C_{n}$ holds almost surely, we immediately derive
$\mathbb{E}[X_{n+1}]\leq(1-\alpha_{n+1})\mathbb{E}[X_{n}]+\mathbb{E}[C_{n}]$
and Lemma 4.1 yields $\mathbb{E}[X_{n}]\leq 2L/(n+2)$. Proceeding as in the proof of Lemma 2.3, noting that we in particular have $X_{n+1}\leq X_{n}+C_{n}$ almost surely, we similarly derive
$\mathbb{P}(\exists n\geq N(U_{n}\geq\varepsilon))\leq\frac{1}{\varepsilon}\left(\mathbb{E}[X_{N}]+\sum_{i=N}^{\infty}\mathbb{E}[C_{i}]\right)$
for $U_{n}:=X_{n}+\sum_{i=n}^{\infty}C_{i}$. In particular, we have $\mathbb{E}[X_{N}]\leq 2L/(N+2)$ and
$\sum_{i=N}^{\infty}\mathbb{E}[C_{i}]\leq L\sum_{i=N}^{\infty}(\alpha_{i}-\alpha_{i+1})=L\alpha_{N}=\frac{2L}{N+2}$
so that
$\mathbb{P}(\exists n\geq N(U_{n}\geq\varepsilon))\leq\frac{1}{\varepsilon}\frac{4L}{N+2}.$
This gives
$\mathbb{P}\left(\exists n\geq N\left(X_{n}\geq\varepsilon\right)\right)\leq\mathbb{P}\left(\exists n\geq N\left(U_{n}\geq\varepsilon\right)\right)\leq\frac{1}{\varepsilon}\frac{4L}{N+2}$
again as in Lemma 2.3. ∎
Remark 4.3.
Note that from the conclusions of Lemma 4.2, it is rather immediate to give corresponding rates for $\mathbb{E}[X_{n}]\to 0$ and $X_{n}\to 0$ almost surely, e.g. by setting $\Phi(\lambda,\varepsilon):=\left\lceil 4L/\varepsilon\lambda\right\rceil$ for the latter, but we prefer the above formulations in this section to make the constants very explicit.
4.2. Linear rates of asymptotic regularity
We now begin by establishing linear rates of asymptotic regularity for the iterations $(x_{n})$ and $(y_{n})$ in the special case of parameters
$\alpha_{n}=\frac{2}{n+2}\text{ and }\beta_{n}=\beta\in(0,1).$
Theorem 4.4.
Let $(x_{n}),(y_{n})$ be the sequences generated by (1) for the parameters as in (4.2). Assume (Hyp) with constant $K_{0}$. Also, assume that $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]\leq K_{1}/(n+2)^{2}$ and $\mathbb{E}[\left\lVert\delta_{n}\right\rVert]\leq K_{2}/(n+2)^{2}$. Then
$\mathbb{E}[\left\lVert x_{n}-x_{n+1}\right\rVert]\leq\frac{2L}{n+2}\text{ and }\mathbb{P}\left(\exists i\geq n\left(\left\lVert x_{i}-x_{i+1}\right\rVert\geq\varepsilon\right)\right)\leq\frac{1}{\varepsilon}\frac{4L}{n+2}$
for all $n\in\mathbb{N}$ where $L=2K_{0}+2K_{1}+2K_{2}$ in both cases.
Proof.
As in the proof of Theorem 3.2, we have $X_{n+1}\leq(1-\alpha_{n+1})X_{n}+C_{n}$ for $X_{n}:=\left\lVert x_{n}-x_{n+1}\right\rVert$ and
$C_{n}:=\left\lVert\xi_{n+1}-\xi_{n}\right\rVert+\left\lVert\delta_{n+1}-\delta_{n}\right\rVert+(\alpha_{n}-\alpha_{n+1})(\left\lVert Tx_{n}-u\right\rVert+\left\lVert\xi_{n}\right\rVert).$
Also following the proof of Theorem 3.2 we have
$\mathbb{E}[X_{0}]=\mathbb{E}[\left\lVert x_{0}-x_{1}\right\rVert]\leq 2K_{0}+K_{1}+K_{2}\leq L.$
So it remains to show that $\mathbb{E}[C_{n}]\leq(\alpha_{n}-\alpha_{n+1})L$, and for this it suffices to show that
$\mathbb{E}[\left\lVert\xi_{n+1}-\xi_{n}\right\rVert]\leq(\alpha_{n}-\alpha_{n+1})\cdot 2K_{1}\text{ and }\mathbb{E}[\left\lVert\delta_{n+1}-\delta_{n}\right\rVert]\leq(\alpha_{n}-\alpha_{n+1})\cdot 2K_{2}.$
We conclude by observing that
$\mathbb{E}[\left\lVert\xi_{n+1}-\xi_{n}\right\rVert]$ $\leq\mathbb{E}[\left\lVert\xi_{n+1}\right\rVert]+\mathbb{E}[\left\lVert\xi_{n}\right\rVert]=\frac{K_{1}}{(n+3)^{2}}+\frac{K_{1}}{(n+2)^{2}}$
$\leq\frac{2K_{1}}{(n+2)^{2}}\leq\frac{4K_{1}}{(n+2)(n+3)}=(\alpha_{n}-\alpha_{n+1})\cdot 2K_{1}$
and similarly for $(\delta_{n})$ and $K_{2}$. The rates then follow from Lemma 4.2. ∎
Remark 4.5.
Before moving to the other asymptotic regularity results, we just briefly note that the asymptotic condition $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]\leq K_{1}/(n+2)^{2}$ naturally implies that $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]<\infty$ with a rather simple rate of convergence that can be easily calculated from the fact that
$\sum_{n=N}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$ $\leq K_{1}\sum_{n=N}^{\infty}\frac{1}{(n+2)^{2}}\leq K_{1}\sum_{n=N}^{\infty}\frac{1}{(n+1)(n+2)}$
$=K_{1}\sum_{n=N}^{\infty}\left(\frac{1}{n+1}-\frac{1}{n+2}\right)=\frac{K_{1}}{N+1}$
for $N\geq 1$. Similarly, this applies to $\delta_{n}$ and $K_{2}$. In particular, as highlighted before in Remark 3.6, we have
$\mathbb{P}(\exists n\geq N(\left\lVert\xi_{n}\right\rVert\geq\varepsilon))\leq\sum_{n=N}^{\infty}\mathbb{P}(\left\lVert\xi_{n}\right\rVert\geq\varepsilon)\leq\sum_{n=N}^{\infty}\frac{\mathbb{E}[\left\lVert\xi_{n}\right\rVert]}{\varepsilon}\leq\frac{1}{\varepsilon}\frac{K_{1}}{N+1}\leq\frac{1}{\varepsilon}\frac{2K_{1}}{N+2}.$
Now, in the case of sequence $(y_{n})$, the above then immediately implies the following:
Theorem 4.6.
Let $(x_{n}),(y_{n})$ be the sequences generated by (1) for the parameters as in (4.2). Assume (Hyp) with constant $K_{0}$. Also, assume that $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]\leq K_{1}/(n+2)^{2}$ and $\mathbb{E}[\left\lVert\delta_{n}\right\rVert]\leq K_{2}/(n+2)^{2}$. Then
$\mathbb{E}[\left\lVert y_{n}-y_{n+1}\right\rVert]\leq\frac{2L}{n+2}$
for all $n\in\mathbb{N}$ where $L$ can be given as an integer linear combination of $K_{0}$, $K_{1}$ and $K_{2}$. If we assume (Hyp) with $K_{0}$ and $Y$, then
$\mathbb{P}\left(\exists i\geq n\left(\left\lVert y_{i}-y_{i+1}\right\rVert\geq\varepsilon\right)\right)\leq\frac{1}{\varepsilon}\frac{4L}{n+2}$
for all $n\in\mathbb{N}$ with a suitable $L$ constructed similarly.
Proof.
Using Lemma 3.1, (1), we have
$\left\lVert y_{n+1}-y_{n}\right\rVert\leq\left\lVert x_{n+1}-x_{n}\right\rVert+\left\lVert\xi_{n+1}-\xi_{n}\right\rVert+\alpha_{n}\left(\left\lVert Tx_{n}-u\right\rVert+\left\lVert\xi_{n}\right\rVert\right)$
for all $n\in\mathbb{N}$ pointwise everywhere. By taking expectations, we get
$\mathbb{E}[\left\lVert y_{n+1}-y_{n}\right\rVert]\leq\mathbb{E}[\left\lVert x_{n+1}-x_{n}\right\rVert]+\mathbb{E}[\left\lVert\xi_{n+1}-\xi_{n}\right\rVert]+\alpha_{n}(K_{0}+K_{1}).$
From Theorem 4.4, we get $\mathbb{E}[\left\lVert x_{n}-x_{n+1}\right\rVert]\leq 2L/(n+2)$ and similar as in the proof thereof, we have $\mathbb{E}[\left\lVert\xi_{n+1}-\xi_{n}\right\rVert]\leq 2K_{1}/(n+2)^{2}\leq 2K_{1}/(n+2)$. Combined with the definition of $\alpha_{n}$, we get the first claim for a suitable $L$ arising as an integer linear combination of $K_{0}$, $K_{1}$ and $K_{2}$. The second claim follows similarly, noting the above Remark 4.5 and the fact that, using Markov’s inequality, we have
$\mathbb{P}\left(\exists n\geq N\left(\alpha_{n}(\left\lVert Tx_{n}-u\right\rVert+\left\lVert\xi_{n}\right\rVert)\geq\varepsilon\right)\right)$
$\qquad\leq\mathbb{P}\left(\exists n\geq N\left((\left\lVert Tx_{n}-u\right\rVert+\left\lVert\xi_{n}\right\rVert)\geq\varepsilon/\alpha_{N}\right)\right)$
$\qquad\leq\mathbb{P}\left(\exists n\geq N\left(Y\geq\varepsilon/2\alpha_{N}\right)\right)+\mathbb{P}\left(\exists n\geq N\left(\left\lVert\xi_{n}\right\rVert\geq\varepsilon/2\alpha_{N}\right)\right)$
$\qquad\leq\frac{1}{\varepsilon}\frac{4K_{0}}{N+2}+\frac{1}{\varepsilon}\frac{8K_{1}}{N+2},$
which, combined with the previous, rather immediately yields the result (which we therefore do not spell out any further). ∎
Remark 4.7.
Reformulated, the above results in particular state that if $(x_{n}),(y_{n})$ are the sequences generated by (1) for parameters as in (4.2) under the assumption (Hyp) and $\mathbb{E}(\left\lVert\xi_{n}\right\rVert)=O(1/n^{2})$ as well as $\mathbb{E}(\left\lVert\delta_{n}\right\rVert)=O(1/n^{2})$, then
$\mathbb{E}[\left\lVert x_{n}-x_{n+1}\right\rVert]=O(1/n)\text{ and }\mathbb{E}[\left\lVert y_{n}-y_{n+1}\right\rVert]=O(1/n).$
4.3. Linear rates of asymptotic regularity relative to the mappings in special cases
In the special case of the stochastic Halpern iteration, which we reobtain (as discussed before) by setting $U:=\mathrm{Id}$ as well as $\delta_{n}:=0$, we get the following fast rates in the above special case:
Theorem 4.8.
Let $(x_{n}),(y_{n})$ be the sequences generated by (1) for parameters as in (4.2) and where $U:=\mathrm{Id}$ and $\delta_{n}:=0$. Assume (Hyp) with constant $K_{0}$. Also, assume that $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]\leq K_{1}/(n+2)^{2}$. Then
$\mathbb{E}[\left\lVert Tx_{n}-x_{n}\right\rVert]\leq\frac{2L}{n+2}$
for all $n\in\mathbb{N}$ where $L$ can be given as an integer linear combination of $K_{0}$ and $K_{1}$. If we assume (Hyp) with $K_{0}$ and $Y$, then
$\mathbb{P}\left(\exists i\geq n\left(\left\lVert Tx_{i}-x_{i}\right\rVert\geq\varepsilon\right)\right)\leq\frac{1}{\varepsilon}\frac{4L}{n+2}$
for all $n\in\mathbb{N}$ with a suitable $L$ constructed similarly.
Proof.
Using the inequalities listed in the proof of Theorem 3.4, we obtain
$\left\lVert Tx_{n}-x_{n}\right\rVert\leq 3\left\lVert x_{n+1}-x_{n}\right\rVert+\alpha_{n}\left\lVert Tx_{n}-u\right\rVert+\left\lVert\xi_{n}\right\rVert$
for all $n\in\mathbb{N}$ pointwise everywhere, in this special case where $U=\mathrm{Id}$ and $\delta_{n}=0$. This immediately yields the above rates (using similar arguments as in Lemma 4.6 in the case of the almost sure convergence) using the previous Theorem 4.4 (noting that in this case $K_{2}=0$). ∎
Theorem 4.8 is closely related to [4, Theorem 3.3], but with adjusted step-sizes that now provide exact linear rates (without logarithmic factors).
In the special case of the stochastic Krasnoselskii-Mann iteration with Tikhonov regularization terms, which we re-obtain by setting $T:=\mathrm{Id}$ as well as $\xi_{n}:=0$, we get the following fast rates in the above special case:
Theorem 4.9.
Let $(x_{n}),(y_{n})$ be the sequences generated by (1) for parameters as in (4.2) and where $T:=\mathrm{Id}$ and $\xi_{n}:=0$. Assume (Hyp) with constant $K_{0}$. Also, assume that $\mathbb{E}[\left\lVert\delta_{n}\right\rVert]\leq K_{2}/(n+2)^{2}$. Lastly, let $B\geq 1/(1-\beta)$. Then
$\mathbb{E}[\left\lVert Ux_{n}-x_{n}\right\rVert]\leq\frac{2L}{n+2}$
for all $n\in\mathbb{N}^{*}$ where $L$ can be constructed in terms of $K_{0}$, $K_{2}$ and $B$. If we assume (Hyp) with $K_{0}$ and $Y$, then
$\mathbb{P}\left(\exists i\geq n\left(\left\lVert Ux_{i}-x_{i}\right\rVert\geq\varepsilon\right)\right)\leq\frac{1}{\varepsilon}\frac{4L}{n+2}$
for all $n\in\mathbb{N}^{*}$ with a suitable $L$ constructed similarly.
Proof.
Using Lemma 3.7, we have
$\left\lVert Uy_{n}-y_{n}\right\rVert\leq B\left(2\left\lVert y_{n-1}-y_{n}\right\rVert+\alpha_{n}\left\lVert Uy_{n}-u\right\rVert+\left\lVert\delta_{n-1}\right\rVert\right)$
for $n\geq 1$ pointwise everywhere. Using the preceding Theorem 4.6, we immediately get that
$\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\leq\frac{2L_{0}}{n+2}\text{ and }\mathbb{P}\left(\exists i\geq n\left(\left\lVert Uy_{i}-y_{i}\right\rVert\geq\varepsilon\right)\right)\leq\frac{1}{\varepsilon}\frac{4L_{0}}{n+2}$
for $n\geq 1$ and a suitable constant $L_{0}$ arising as an integer linear combination of $K_{0}$ and $K_{2}$. Using the inequalities from Theorem 3.4, we then further have
$\left\lVert Ux_{n}-x_{n}\right\rVert\leq 2\left\lVert x_{n+1}-x_{n}\right\rVert+3\left\lVert Uy_{n}-y_{n}\right\rVert+2\left\lVert\delta_{n}\right\rVert$
for all $n\in\mathbb{N}$ pointwise everywhere and so, using the previous results as well as Theorem 4.4, we get the desired rates. ∎
4.4. Fast rates of asymptotic regularity relative to the mappings in the general case
In the context of the above assumptions on the scalar sequences and the errors, we can still get rather sensible complexity estimates in the general case where neither mapping necessarily trivializes. While we could express this again using a general modulus of uniform convexity $\eta$ for the underlying space, we here focus on the case where $\eta$ is of power type $p$ for $p\geq 2$, i.e. there exists a constant $C$ such that $\eta(\varepsilon)=C\varepsilon^{p}$. Crucially, this is the case for inner product spaces for $p=2$:
Lemma 4.10 (essentially [25]).
If $X$ is a inner product space, then $X$ is uniformly convex with a corresponding modulus $\eta(\varepsilon)=\varepsilon^{2}/8$.
This allows for the following results on the asymptotic regularity in the general case. We begin by instantiating Theorem 3.12 on the asymptotic regularity of the sequence $(y_{n})$ relative to $U$ almost surely and Theorem 3.18 for deriving the respective regularity result in expectation.
Lemma 4.11.
Let $X$ be uniformly convex with a modulus $\eta$ of power type $p$ with constant $C$. Let $(x_{n}),(y_{n})$ be the sequences generated by (1) for parameters as in (4.2). Let $K$ and $Y$ be as in Lemma 3.9. Also, assume that $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]\leq K_{1}/(n+2)^{2}$ and $\mathbb{E}[\left\lVert\delta_{n}\right\rVert]\leq K_{2}/(n+2)^{2}$. Lastly, let $\Lambda>0$ be such that $\Lambda\leq\beta\leq 1-\Lambda$. Then $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely with rate
$\Gamma(\lambda,\varepsilon):=\left\lceil\frac{(3K)^{p-1}L}{C\Lambda^{2}\varepsilon^{p}\lambda^{p}}\right\rceil$
for a suitable $L$ arising as an integer linear combination of $K$, $K_{1}$ and $K_{2}$. Assuming the existence of a modulus $\mu$ of uniform integrability for $(\left\lVert Uy_{n}-y_{n}\right\rVert)$, we further get $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ with rate
$\Gamma^{\prime}(\varepsilon):=\left\lceil\frac{2^{p}(3K)^{p-1}L}{C\Lambda^{2}\varepsilon^{p}\mu(\varepsilon/4)^{p}}\right\rceil$
where $L$ is as above.
Proof.
First note that in the context of moduli $\eta$ of power type $p\geq 2$, we are actually in the setting of the previous Remarks 3.13 and 3.19 where $\tilde{\eta}(\varepsilon)=C\varepsilon^{p-1}$. Then the rate for $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely follows by instantiating the rate given in Theorem 3.12 with the following moduli: With the above $\tilde{\eta}$, we have $\widehat{\varepsilon}:=C\Lambda^{2}\varepsilon^{p}\lambda^{p-1}/(3K)^{p-1}$ and $K^{\prime}:=3K/\lambda$. Using Theorem 4.4, we have
$\mathbb{P}\left(\exists i\geq n\left(\left\lVert x_{i}-x_{i+1}\right\rVert\geq\varepsilon\right)\right)\leq\frac{1}{\varepsilon}\frac{4L_{0}}{n+2}$
for all $n\in\mathbb{N}$ and a suitable constant $L_{0}$ arising as an integer linear combination of $K$, $K_{1}$ and $K_{2}$. So we in particular have that $\Delta(\lambda,\varepsilon)=\left\lceil 4L_{0}/\varepsilon\lambda\right\rceil$ is a corresponding rate for $\left\lVert x_{n}-x_{n+1}\right\rVert\to 0$ almost surely. As $\alpha=2/(n+2)$, we further have rather immediately that $\rho(\varepsilon)=\left\lceil 2/\varepsilon\right\rceil$ is a corresponding rate for $\alpha_{n}\to 0$. Lastly, using the assumptions on $\left\lVert\xi_{n}\right\rVert$ and $\left\lVert\delta_{n}\right\rVert$, note that as in Remark 4.5 we have $\sum_{n=N}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]\leq K_{1}/(N+1)$ so that $\chi_{1}(\varepsilon)=\left\lceil K_{1}/\varepsilon\right\rceil$ is a corresponding rate of convergence for $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]<\infty$. The rate of convergence $\chi_{2}(\varepsilon)=\left\lceil K_{2}/\varepsilon\right\rceil$ for $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<\infty$ follows similarly. Then instantiating Theorem 3.12 under Remark 3.13 gives us the rate
$\max\left\{\left\lceil\frac{144L_{0}(3K)^{p-1}}{C\Lambda^{2}\varepsilon^{p}\lambda^{p}}\right\rceil,\left\lceil\frac{24K(3K)^{p-1}}{C\Lambda^{2}\varepsilon^{p}\lambda^{p}}\right\rceil,\left\lceil\frac{36K_{1}(3K)^{p-1}}{C\Lambda^{2}\varepsilon^{p}\lambda^{p}}\right\rceil,\left\lceil\frac{36K_{2}(3K)^{p-1}}{C\Lambda^{2}\varepsilon^{p}\lambda^{p}}\right\rceil\right\}\leq\frac{(3K)^{p-1}L}{C\Lambda^{2}\varepsilon^{p}\lambda^{p}}$
for $L:=144L_{0}$, noting that $K,K_{1},K_{2}\leq L_{0}$ and so the first part follows. For the second part, we just apply Lemma 3.17 directly. ∎
Using that, we can then employ the previous Theorems 3.4 and 3.5 to derive rates of asymptotic regularity also for the sequence $(x_{n})$ relative to the mappings $U$ and $T$. For simplicity, we now focus on inner product spaces, i.e. where $p=2$ and $C=1/8$ by Lemma 4.10.
Theorem 4.12.
Let $X$ be an inner product space. Let $(x_{n}),(y_{n})$ be the sequences generated by (1) for parameters as in (4.2). Let $K$ and $Y$ be as in Lemma 3.9. Also, assume that $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]\leq K_{1}/(n+2)^{2}$ and $\mathbb{E}[\left\lVert\delta_{n}\right\rVert]\leq K_{2}/(n+2)^{2}$. Lastly, let $\Lambda>0$ be such that $\Lambda\leq\beta\leq 1-\Lambda$. Then $\left\lVert Ux_{n}-x_{n}\right\rVert\to 0$ and $\left\lVert Tx_{n}-x_{n}\right\rVert\to 0$ almost surely with rates
$\Phi_{1}(\lambda,\varepsilon):=\left\lceil\frac{24KL}{\Lambda^{2}\varepsilon^{2}\lambda^{2}}\right\rceil\text{ and }\Phi_{2}(\lambda,\varepsilon):=\left\lceil\frac{72KL}{\Lambda^{2}\varepsilon^{2}\lambda^{2}}\right\rceil,$
respectively, where $L$ is as in Lemma 4.11. Assuming the existence of a modulus $\mu$ of uniform integrability for $(\left\lVert Uy_{n}-y_{n}\right\rVert)$, we further get $\mathbb{E}[\left\lVert Ux_{n}-x_{n}\right\rVert]\to 0$ and $\mathbb{E}[\left\lVert Tx_{n}-x_{n}\right\rVert]\to 0$ with respective rates
$\varphi_{1}(\varepsilon):=\left\lceil\frac{96KL}{\Lambda^{2}\varepsilon^{2}\mu(\varepsilon/4)^{2}}\right\rceil\text{ and }\varphi_{2}(\varepsilon):=\left\lceil\frac{288KL}{\Lambda^{2}\varepsilon^{2}\mu(\varepsilon/4)^{2}}\right\rceil.$
Proof.
The rates follow immediately by instantiating Theorems 3.4 and 3.5 with the rates obtained from Theorem 4.4 and Lemma 4.11, noting in particular that the quadratic rates for $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ and $\mathbb{E}[\left\lVert Uy_{n}-y_{n}\right\rVert]\to 0$ dominate. ∎
5. Further remarks
We conclude the paper with an extended discussion on two further topics: oracle complexities and applications in reinforcement learning. While both of these topics are being explored in greater depth in ongoing future work, we want to sketch key aspects of these developments here.
5.1. Abstract oracle complexities
The intuition behind our abstract stochastic scheme is that $\xi_{n}$ and $\delta_{n}$ arise by evaluating stochastic oracles $\tilde{T}$ and $\tilde{U}$ for $T$ and $U$ respectively. Though our primary focus has been on establishing direct convergence rates for our schemes, in concrete applications the resulting oracle complexity may act as a more reliable estimate of the actual cost of running the algorithm. To address this at a level of generality in line with our overall approach, we propose an abstract characterisation of oracle complexity. We first introducing a pair of sequences of natural numbers (#$\#\xi_{n}$) and (#$\#\delta_{n}$), with the intuition that #$\#\xi_{n}$ denotes the number of calls made to the stochastic oracle $\tilde{T}$ in computing the approximation to $Tx_{n}$ (with output $Tx_{n}+\xi_{n}$), and #$\#\delta_{n}$ the number of calls to $\tilde{U}$ in computing the approximation to $Uy_{n}$.
With this intuition in mind, we say that $\psi:\mathbb{N}\to\mathbb{N}$ is a bound on the oracle growth of the general scheme (1) if
$\forall N\in\mathbb{N}\,\left(\#\xi_{N}+\sum_{n=0}^{N-1}(\#\xi_{n}+\#\delta_{n})\leq\psi(N)\right)$
so that the total number of oracle calls needed to compute $x_{N}$ and $y_{N}$ is bounded by $\psi(N)$. Then, for example, if
$\mathbb{E}[\left\lVert Tx_{n}-x_{n}\right\rVert]\to 0$ with rate $\varphi$,
we can compute for any $\varepsilon>0$ an approximant $x_{N}$ such that
$\mathbb{E}[\left\lVert Tx_{N}-x_{N}\right\rVert]<\varepsilon$ using at most $(\psi\circ\varphi)(\varepsilon)$ queries to the stochastic oracles.
Indeed, the bounding function $\psi$ can be used in a similar way to convert any of our convergence results to a corresponding characterisation of the overall oracle complexity in a completely general manner. One possible instantiation of this abstract perspective is discussed in the following example:
Example 5.1.
Assume that $\tilde{T}$ is evaluated via a minibatch strategy (as e.g. studied in [4]), which would correspond to setting
$\xi_{n}:=\frac{1}{k_{n}}\sum_{j=1}^{k_{n}}\tilde{T}(x_{n},\zeta_{n,j})-Tx_{n}$
for batchsizes $(k_{n})$ (where $\zeta_{n,1},\ldots,\zeta_{n,k_{n}}$ are i.i.d. random variables sampled from a distribution associated with $x_{n}$) and therefore #$\#\xi_{n}=k_{n}$. Under suitable additional variance assumptions on the oracle $\tilde{T}$, one can then show that $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]\leq c/\sqrt{k_{n}}$ for some corresponding constant $c>0$. Hence, in such situations, the oracle bounds then involve a payoff between the rate of variance reduction and the oracle growth, or more concretely between the
($\ast$) convergence speed of 1/kn0 and growth of $\text{convergence speed of $1/\sqrt{k_{n}}\to 0$ and growth of $\sum_{n=0}^{N}k_{n}$}.$
One concrete such situation is e.g. discussed in [4, Corollary 3.5] where, in the context of a finite dimensional space $X$, the oracle $\tilde{T}$ is assumed to have uniformly bounded variance in the sense that
$\sup_{x\in X}\mathbb{E}\left[\left\lVert\tilde{T}(x,\zeta)-Tx\right\rVert^{2}_{2}\right]\leq\sigma^{2}$
[…]
In the special cases of the stochastic Halpern iteration ($U:=\mathrm{Id}$ and $\delta_{n}:=0$) or the stochastic Krasnoselskii-Mann iteration with Tikhonov regularization terms ($T:=\mathrm{Id}$ and $\xi_{n}:=0$), under the additional conditions (Hyp) with constant $K_{0}$ on the boundedness and (4.2) on the parameters as in Theorems 4.8 and 4.9, said theorems immediately imply
$\mathbb{E}[\left\lVert Tx_{n}-x_{n}\right\rVert]\leq\frac{2L}{n+2}\text{ and }\mathbb{E}[\left\lVert Ux_{n}-x_{n}\right\rVert]\leq\frac{2L^{\prime}}{n+2}$
for suitable $L$ and $L^{\prime}$ constructed from $K_{0},K_{1}$ or $K_{0},K_{2}$ and $B\geq 1/(1-\beta)$, respectively. These estimates correspond to rates of convergence $\varphi(\varepsilon):=\varepsilon/2L$ and $\varphi^{\prime}(\varepsilon):=\varepsilon/2L^{\prime}$, respectively, and therefore, in that context, we can compute an $x_{N}$ such that
$\mathbb{E}[\left\lVert Tx_{N}-x_{N}\right\rVert]<\varepsilon\text{ and }\mathbb{E}[\left\lVert Ux_{N}-x_{N}\right\rVert]<\varepsilon$
using at most $\left(\varepsilon/2L\right)^{5}$ or respectively $\left(\varepsilon/2L^{\prime}\right)^{5}$ queries to the stochastic oracles $\tilde{T}$ or $\tilde{U}$ in question. In this way, we in particular obtain a version of [4, Corollary 3.5] without logarithmic factors.
As is to be expected when augmenting existing deterministic methods with noise terms, all our convergence results require us to control the means of $(\xi_{n})$ and $(\delta_{n})$ in some way, and this is intuitively achieved in concrete settings by increasing the number of oracle queries (#$(\#\xi_{n})$ and (#$(\#\delta_{n})$, so that, as above with ($\ast$5.1), obtaining good bounds on the oracle complexity for instances of (1) will generally involve a tension between variance control and oracle growth. Because our explicit rates of convergence make completely precise how improved control on the means through e.g. faster convergence rates for $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$ and $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]$ leads to improved rates of convergence for the algorithm itself, we anticipate that in any given model that allows us to explicitly describe $\psi$ in terms of $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$ and $\mathbb{E}[\left\lVert\delta_{n}\right\rVert]$, our abstract quantitative results would not only provide us with a general bound on the associated oracle complexity in terms of parameters representing variance control, but might even provide insights into specific choices of parameters that optimize oracle complexity.
5.2. Applications in reinforcement learning
Our new stochastic iteration scheme (1) immediately leads to new algorithms in relevant areas of application, including reinforcement learning, where in particular noisy methods for computing fixpoints underlie model-free learning algorithms such as $Q$-learning [39]. Here, direct access to the corresponding operators $T,U$ would require full knowledge of the agent’s environment, so these are instead sampled, leading to iteration schemes with noise terms. Only very recently in [4] has the special case of the scheme (1) corresponding to Halpern’s iteration (obtained by setting $U:=\text{Id}$ and $\delta_{n}:=0$ in (1)) been instantiated to yield a Halpern-type version of $Q$-learning, and in a similar way our general method represents an expanded class of learning algorithms.
To illustrate this, suppose that $(S,A,r,p)$ forms a Markov decision process (MDP) over some finite set of states $S$ and actions $A$, where if we choose action $a$ in state $s$, $r(s,a)$ represents an immediate reward and $p(s,a,t)$ the probability that we transition to state $t$. Concretely, by appropriately instantiating our stochastic Krasnoselskii-Mann iteration with Tikhonov regularization terms (obtained by setting $T:=\text{Id}$ and $\xi_{n}:=0$ as well as $\gamma_{n}:=1-\alpha_{n}$ and $u:=0$ in (1)), we are lead to the following novel $Q$-learning method with Tikhonov regularization terms:
(KM-T-Q) $Q_{n+1}(s,a):=(1-\beta_{n})\left(U\left(\gamma_{n}Q_{n}\right)(s,a)+\delta_{n}(s,a)\right)+\beta_{n}\left(\gamma_{n}Q_{n}(s,a)\right)$
for
$(UQ)(s,a):=r(s,a)+\sum_{t\in S}p(s,a,t)\max_{b\in A}Q(t,b)+c$
for some constant $c\in\mathbb{R}$, where here $U$ is clearly nonexpansive with respect to
$\left\lVert Q\right\rVert_{\infty}=\max_{s\in\mathcal{S}}\max_{a\in\mathcal{A}}Q(s,a).$
There are standard conditions on our MDP (see e.g. [35]) under which an optimal average reward, i.e. a maximum $v^{\ast}$ for
$v_{\pi}:=\liminf_{N\to\infty}\mathbb{E}\left[\frac{1}{N}\sum_{n=0}^{N-1}r(s_{n},a_{n})\right]\text{with $(s_{n},a_{n})$ induced by the policy $\pi$}$
exists and is independent of the initial state, and moreover, an optimal policy obtaining this maximum can be defined as
$\pi^{\ast}(s):=\operatorname*{arg\,max}_{a\in A}Q^{\ast}(s,a)$
where $Q^{\ast}$ is a fixpoint of $U$ for $c:=-v^{\ast}$. Thus the convergence of (KM-T-Q) and similar schemes towards fixpoints is of direct practical relevance in computing optimal policies. In particular, instantiating the noise terms according to a regularized minibatch strategy as e.g. discussed in [4] for a Halpern-type version of $Q$-learning, i.e.
$\delta_{n}(s,a):=\gamma_{n}\left(\frac{1}{s_{n}}\sum_{j=1}^{s_{n}}\max_{b\in S}Q_{n}(\zeta_{n,j},b)-\sum_{t\in S}p(s,a,t)\max_{b\in A}Q_{n}(t,b)\right)$
where $\zeta_{n,1}\ldots,\zeta_{n,k_{n}}$ are i.i.d. random variables sampled from $p(s,a,\cdot)$, (KM-T-Q) then becomes
$Q_{n+1}(s,a):=\gamma_{n}\left((1-\beta_{n})\left(r(s,a)+\frac{1}{s_{n}}\sum_{j=1}^{s_{n}}\max_{b\in S}Q_{n}(\zeta_{n,j},b)-v^{\ast}\right)+\beta_{n}Q_{n}(s,a)\right)$
the convergence of which can then be used (as in [4]) to establish the convergence of concrete variants of RVI-$Q$-learning in the style of the Krasnoselskii-Mann iteration with Tikhonov regularization terms, where in the above iteration $v^{\ast}$ is replaced by some approximant $f(Q_{n})$ for a suitable function $f$.
However, our focus in this paper is not the development of concrete learnings algorithms, but the establishment of abstract convergence results that are of general relevance across areas of application. In that spirit, we simply observe for now that convergence of the abstract $Q$-learning scheme (KM-T-Q) follows immediately from our main convergence results and inherits the associated rates. For example, in the case of fast rates we obtain the following:
Corollary 5.2.
Let $(Q_{n})$ be the sequence generated by (KM-T-Q) for the case of parameters
$\gamma_{n}=1-\frac{2}{n+2}\text{ and }\beta_{n}=\beta\in(0,1).$
Assume that $\left\lVert Q_{n}\right\rVert_{\infty},\left\lVert r+c\right\rVert_{\infty}\leq K$ so that (Hyp) holds with constant $K_{0}:=2K$. Also, assume that $\mathbb{E}[\left\lVert\delta_{n}\right\rVert_{\infty}]\leq K_{2}/(n+2)^{2}$. Lastly, let $B\geq 1/(1-\beta)$. Then
$\mathbb{E}[\left\lVert UQ_{n}-Q_{n}\right\rVert_{\infty}]\leq\frac{2L}{n+2}$
for all $n\in\mathbb{N}^{*}$ where $L$ is constructed from $K_{0},K_{2}$ and $B$ as in Theorem 4.9.
Remark 5.3.
Instantiating $\delta_{n}$ as above via a regularized minibatch strategy, under conditions as described in Example 5.1 we would obtain the desired control of the means by defining our batchsizes through $s_{n}:=(n+4)^{4}$, which akin to the discussion in Example 5.1 also immediately would allow us an estimation of the oracle complexity in that case which is of pentic complexity.
It would be very interesting to see in future work whether $Q$-learning could, by means of Tikhonov regularization terms as in the scheme (KM-T-Q), benefit from the nice theoretical and practical properties that these regularization terms induce in the non-stochastic case.
We conclude by remarking that much broader generalisations which more fully exploit the convergence results of this paper are possible. One obvious avenue for further study would be to consider alternating $Q$-learning algorithms based on our main scheme (1) and we want to propose on such example by considering the following scheme which represents a Halpern-Mann-type variant of $Q$-learning (where the anchor point is taken to be zero):
(sHM-Q) $Q^{\prime}_{n}(s,a):=(1-\alpha_{n})\left(r(s,a)+\sum_{t\in S}p(s,a,t)\max_{b\in A}Q_{n}(t,b)+\xi_{n}(s,a)\right)$
$Q_{n+1}(s,a):=(1-\beta_{n})\left(r^{\prime}(s,a)+\sum_{t\in S}p^{\prime}(s,a,t)\max_{b\in A}Q^{\prime}_{n}(t,b)+\delta_{n}(s,a)\right)+\beta_{n}Q^{\prime}_{n}$
where $(S,A,r^{\prime},p^{\prime})$ is a second MDP over the same states and actions. This method bears a passing resemblance to double $Q$-learning [16], which incorporates a double estimator to reduce bias in ordinary $Q$-learning. Moreover, our scheme converges to a simultaneous fixpoint of two underlying nonexpansive operators, and thus might be relevant in situations where we are required to compute optimal policies concurrently across distinct environments. However, we leave an exploration of the potential merits of both schemes (KM-T-Q) and (sHM-Q) as well as similar algorithms to future work.
Acknowledgments: The first author was supported by the ‘Deutsche Forschungsgemeinschaft’ Project DFG KO 1737/6-2. The second author was supported by the EPSRC grant EP/W035847/1.